### v1

# Graph Transformers: A Survey

Abstract
Graph transformers are a recent advancement in machine learning, offering a new class of neural network models for graph-structured data. The synergy between transformers and graph learning demonstrates strong performance and versatility across various graph-related tasks. This survey provides an in-depth review of recent progress and challenges in graph transformer research. We begin with foundational concepts of graphs and transformers. We then explore design perspectives of graph transformers, focusing on how they integrate graph inductive biases and graph attention mechanisms into the transformer architecture. Furthermore, we propose a taxonomy classifying graph transformers based on depth, scalability, and pre-training strategies, summarizing key principles for effective development of graph transformer models. Beyond technical analysis, we discuss the applications of graph transformer models for node-level, edge-level, and graph-level tasks, exploring their potential in other application scenarios as well. Finally, we identify remaining challenges in the field, such as scalability and efficiency, generalization and robustness, interpretability and explainability, dynamic and complex graphs, as well as data quality and diversity, charting future directions for graph transformer research.
Index Terms:
Graph transformer, attention, graph neural network, representation learning, graph learning, network embedding
I Introduction
Graphs, as data structures with high expressiveness, are widely used to present complex data in diverse domains, such as social media, knowledge graphs, biology, chemistry, and transportation networks [1]. They capture both structural and semantic information from data, facilitating various tasks, such as recommendation [2], question answering [3], anomaly detection [4], sentiment analysis [5], text generation [6], and information retrieval [7]. To effectively deal with graph-structured data, researchers have developed various graph learning models, such as graph neural networks (GNNs), learning meaningful representations of nodes, edges and graphs [8]. Particularly, GNNs following the message-passing framework iteratively aggregate neighboring information and update node representations, leading to impressive performance on various graph-based tasks [9]. Applications ranging from information extraction to recommender systems have benefited from GNN modelling of knowledge graphs [10].
More recently, the graph transformer, as a newly arisen and potent graph learning method, has attracted great attention in both academic and industrial communities [11, 12]. Graph transformer research is inspired by the success of transformers in natural language processing (NLP) [13] and computer vision (CV) [14], coupled with the demonstrated value of GNNs. Graph transformers incorporate graph inductive bias (e.g., prior knowledge or assumptions about graph properties) to effectively process graph data [15]. Furthermore, they can adapt to dynamic and heterogeneous graphs, leveraging both node and edge features and attributes. [16]. Various adaptations and expansions of graph transformers have shown their superiority in tackling diverse challenges of graph learning, such as large-scale graph processing [17]. Furthermore, graph transformers have been successfully employed in various domains and applications, demonstrating their effectiveness and versatility.
Existing surveys do not adequately cover the latest advancements and comprehensive applications of graph transformers. In addition, most do not provide a systematic taxonomy of graph transformer models. For instance, Chen et al.  [18] focused primarily on the utilization of GNNs and graph transformers in CV, but they failed to summarize the taxonomy of graph transformer models and ignored other domains, such as NLP. Similarly, Müller et al. [12] offered an overview of graph transformers and their theoretical properties, but they did not provide a comprehensive review of existing methods or evaluate their performance on various tasks. Lastly, Min et al. [19] concentrated on the architectural design aspect of graph transformers, offering a systematic evaluation of different components on different graph benchmarks, but they did not include significant applications of graph transformers or discuss open issues in this field.
To fill these gaps, this survey aims to present a comprehensive and systematic review of recent advancements and challenges in graph transformer research from both design and application perspectives. In comparison to existing surveys, our main contributions are as follows:
  1. 1.
    We provide a comprehensive review of the design perspectives of graph transformers, including graph inductive bias and graph attention mechanisms. We classify these techniques into different types and discuss their advantages and limitations.
  2. 2.
    We present a novel taxonomy of graph transformers based on their depth, scalability, and pre-training strategy. We also provide a guide to choosing effective graph transformer architectures for different tasks and scenarios.
  3. 3.
    We review the application perspectives of graph transformers in various graph learning tasks, as well as the application scenarios in other domains, such as NLP and CV tasks.
  4. 4.
    We identify the crucial open issues and future directions of graph transformer research, such as the scalability, generalization, interpretability, and explainability of models, efficient temporal graph learning, and data-related issues.
An overview of this paper is depicted in Figure 1. The subsequent survey is structured as follows: Section II introduces notations and preliminaries pertaining to graphs and transformers. Section III delves into the design perspectives of graph transformers that encompass graph inductive bias and graph attention mechanisms. Section IV presents a taxonomy of graph transformers categorizing them based on their depth, scalability and pre-training strategy. Additionally, a guide is provided for selecting appropriate graph transformer models for diverse tasks and domains. Section V explores the application perspectives of graph transformers on various node-level, edge-level, and graph-level tasks, along with other application scenarios. Section VI identifies open issues and future directions for research on graph transformers. Lastly, Section VII concludes the paper and highlights its main contributions.
Figure 1: Organization of this paper
II Notations and Preliminaries
In this section, we present fundamental notations and concepts utilized throughout this survey paper. Additionally, we provide a concise summary of current methods for graph learning and self-attention mechanisms which serve as the basis for graph transformers. Table I includes mathematical notations used in this paper.
TABLE I: Notations in this paper
Notation Definition
$G=(V,E)$ A graph with node set $V$ and edge set $E$
$N$ Number of nodes
$M$ Number of edges
$\textbf{A}\in\mathbb{R}^{N\times N}$ Adjacency matrix of graph $G$
$\textbf{X}\in\mathbb{R}^{N\times d_{n}}$ Node feature matrix, $\textbf{x}_{i}\in\textbf{X}$
$\textbf{F}\in\mathbb{R}^{M\times d_{e}}$ Edge feature matrix
$\textbf{h}_{v}^{(l)}$ Hidden state of node $v$ at layer $l$
$\phi$ An update function for node states
$\oplus$ An aggregation function for neighbor states
$N(v)$ Neighbor set of node $v$
$f$ A message function for the node and edge states
[…]
$d_{k}$ Dimension of query and key matrices
$\textbf{p}_{i}$ Positional encoding of node $v_{i}$
$d(i,j)$ The shortest path distance between node $v_{i}$ and node $v_{j}$
$\textbf{e}_{ij}$ Edge feature between node $v_{i}$ and node $v_{j}$
$a_{ij}$ Attention score between node $v_{i}$ and node $v_{j}$
$W,b$ Learnable parameters for the self-attention layer
$\textbf{L}=\textbf{I}_{N}-\textbf{D}^{-1/2}\textbf{AD}^{-1/2}$ Normalized graph Laplacian matrix
U Eigenvectors matrix of L
II-A Graphs and Graph Neural Networks
A graph is a data structure consisting of a set of nodes (or vertices) $V$ and a set of edges (or links) $E$ that connect pairs of nodes. Formally, a graph can be defined as $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is node set with $N$ nodes and $E=\{e_{1},e_{2},\dots,e_{M}\}$ is edge set with $M$ edges. Edge $e_{k}=(v_{i},v_{j})$ indicates the connection between node $v_{i}$ and node $v_{j}$, where $i,j\in\{1,2,\dots,N\}$ and $k\in\{1,2,\dots,M\}$. A graph can be represented by an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}$ indicates the presence or absence of an edge between node $v_{i}$ and node $v_{j}$. Alternatively, a graph can be represented by the edge list $E\in\mathbb{R}^{M\times 2}$, where each row of $E$ contains the indices of two nodes connected by an edge. A graph can also have node features and edge features that describe the attributes or properties of nodes and edges, respectively. The features of the nodes can be represented by a feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{n}}$, where $d_{n}$ is the dimension of the node features. The edge features can be represented by a feature tensor $\textbf{F}\in\mathbb{R}^{M\times d_{e}}$, where $d_{e}$ is the dimension of the edge features [20].
Graph learning refers to the task of acquiring low-dimensional vector representations, also known as embeddings for nodes, edges, or the entire graph. These embeddings are designed to capture both structural and semantic information of the graph. GNNs are a type of neural network model that excels at learning from graph-structured data. They achieve this by propagating information along edges and aggregating information from neighboring nodes [21]. GNNs can be categorized into two main groups: spectral methods and spatial methods.
Spectral methods are based on graph signal processing and graph Fourier transform, implementing convolution operations on graphs in the spectral domain [22]. The Fourier graph transform is defined as $\mathbf{\hat{X}}=\textbf{U}^{T}\textbf{X}\textbf{U}$, where $\mathbf{\hat{X}}$ is the spectral representation of node feature matrix X and U is the eigenvector matrix of the normalized graph Laplacian matrix $\textbf{L}=\textbf{I}_{N}-\textbf{D}^{-1/2}\textbf{A}\textbf{D}^{-1/2}$, where $\textbf{I}_{N}$ is the identity matrix and D is diagonal degree matrix with $D_{ii}=\sum_{j=1}^{N}A_{ij}$. Spectral methods can capture global information about the graph, but they suffer from high computational complexity, poor scalability, and the lack of generalization to unseen graphs [23].
Spatial methods are based on message-passing and neighborhood aggregation, implementing convolution operations on graphs in the spatial domain [24]. The message-passing framework is defined as:
$\textbf{h}_{v}^{(l+1)}=\phi\left(\textbf{h}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}f(\textbf{h}_{u}^{(l)},\textbf{h}_{v}^{(l)},\textbf{e}_{uv})\right)\end{aligned},$ (1)
where $\textbf{h}_{v}^{(l)}$ is the hidden state of node $v$ at layer $l$, $\phi$ is an update function and $\oplus$ is an aggregation function. $\mathcal{N}(v)$ is the set of neighbors of node $v$ and $f$ is a message function that depends on node states and edge features. $\textbf{e}_{uv}$ is the feature vector of the edge between nodes $u$ and $v$. Spatial methods can capture local information of the graph, but they have limitations in modeling long-range dependencies, complex interactions and heterogeneous structures [25].
II-B Self-attention and transformers
Self-attention is a mechanism that enables a model to learn to focus on pertinent sections of input or output sequences [26]. It calculates a weighted sum of all elements in a sequence with weights determined by the similarity between each element and a query vector. Formally, self-attention is defined as:
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}\right)\textbf{V}\end{aligned},$ (2)
where Q, K and V are query, key and value matrices,respectively. $d_{k}$ is the dimension of the query and key matrices. Self-attention can capture long-range dependencies, global context and variable-length sequences without using recurrence or convolution.
Figure 2: An illustration of transformer architecture [13].
Transformers are neural network models that use self-attention as the main building block [13]. Transformers consist of two main components: an encoder and a decoder. The encoder takes an input sequence $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\dots,\textbf{x}_{N}\}$ and generates a sequence of hidden states $\textbf{Z}=\{\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{N}\}$. The decoder takes an output sequence $\textbf{Y}=\{\textbf{y}_{1},\textbf{y}_{2},\dots,\textbf{y}_{N}\}$ and generates a sequence of hidden states $\textbf{S}=\{\textbf{s}_{1},\textbf{s}_{2},\dots,\textbf{s}_{N}\}$. The decoder also uses an attention mechanism to attend to the encoder's hidden states. Formally, the encoder and decoder are defined as:
$\textbf{z}_{i}$ $=\text{EncoderLayer}(\textbf{x}_{i},\textbf{Z}_{<i}),$ (3)
$\textbf{s}_{j}$ $=\text{DecoderLayer}(\textbf{y}_{j},\textbf{S}_{<j},\textbf{Z}).$
Here, EncoderLayer and DecoderLayer are composed of multiple self-attention and feed-forward sublayers. Transformers can achieve state-of-the-art results on various tasks, such as machine translation [27], text mining [28, 29], document comprehending [30, 31], image retrieval [32], visual question answering [3], image generation [33], and recommendation systems [34, 35]. An overview of vanilla transformer is shown in Figure 2.
Graph transformers integrate graph inductive bias into transformer to acquire knowledge from graph-structured data [36]. By employing self-attention mechanisms on nodes and edges, graph transformers can effectively capture both local and global information of the graph. In particular, graph transformers exhibit the ability to handle heterogeneous graphs containing diverse types of nodes and edges, as well as complex graphs featuring higher-order structures [12, 37].
III Design Perspectives of graph transformers
In this section, we discuss the primary architectures of graph transformers, aiming to explore their design perspectives in depth. Particularly, we will focus on two key components: graph inductive biases and graph attention mechanisms, to understand how these elements shape graph transformer models’ capabilities.
III-A Graph Inductive Bias
Unlike Euclidean data, such as texts and images, graph data is non-Euclidean data, which has intricate structures and lacks a fixed order and dimensionality, posing difficulties in directly applying standard transformers on graph data [16]. To address this issue, graph transformers incorporate graph inductive bias to encode the structural information of graphs and achieve effective generalization of transformers across new tasks and domains. In this section, we explore the design perspectives of graph transformers through the lens of graph inductive bias. We classify graph inductive bias into four categories: node positional bias, edge structural bias, message-passing bias, and attention bias (see Figure. 3).
Figure 3: Graph Inductive bias: (A) and (B) are local and global node positional biases, respectively. (C) and (D) are local and global edge structural biases, respectively. (E) and (F) are local and global attention biases, respectively. Notably, message-passing bias is a kind of local attention bias.
III-A1 Node Positional Bias
Node positional bias is a crucial inductive bias for graph transformers because it provides information about the relative or absolute positions of nodes in a graph [38]. Formally, given a graph $G=(V,E)$ with $N$ nodes and $M$ edges, each node $v_{i}\in V$ has a feature vector $\textbf{x}_{i}\in\mathbb{R}^{d}_{n}$. A graph transformer aims to learn a new feature vector $\textbf{h}_{i}\in\mathbb{R}^{d}_{k}$ for each node by applying a series of self-attention layers. A self-attention layer can be defined as:
$\textbf{h}_{i}=\sum_{j=1}^{n}a_{ij}W\textbf{x}_{j}+b,$ (4)
where $a_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, measureing the relevance or similarity of their features. $W$ and $b$ are learnable parameters. However this self-attention mechanism does not consider the structural and positional information of nodes, which is crucial for capturing graph semantics and inductive biases [39]. Node positional encoding is a way to address this challenge by providing additional positional features to nodes, reflecting their relative or absolute positions in the graph.
Local Node Positional Encodings. Building on the success of relative positional encodings in NLP, graph transformers leverage a similar concept for local node positional encodings. In NLP, each token receives a feature vector that captures its relative position and relationship to other words in the sequence [40]. Likewise, graph transformers assign feature vectors to nodes based on their distance and relationships with other nodes in the graph [41]. This encoding technique aims to preserve the local connectivity and neighborhood information of nodes, which is critical for tasks like node classification, link prediction, and graph generation.
A proficient approach for integrating local node positional information involves the utilization of one-hot vectors. These vectors represent the hop distance between a node and its neighboring nodes [11]:
$\textbf{p}_{i}=[I(d(i,j)=1),I(d(i,j)=2),...,I(d(i,j)=max)].$ (5)
In this equation, $d(i,j)$ represents the shortest path distance between node $v_{i}$ and node $v_{j}$ and $I$ is an indicator function that returns 1 if its argument is true and 0 otherwise. The maximum hop distance is denoted by $max$. This encoding technique was utilized by Velickovic et al. [42] to enhance their Graph Attention Networks (GATs) with relative position-aware self-attention. Another way to incorporate local node positional encodings in a graph is by using learnable embeddings that capture the relationship between two nodes [43, 44]. This approach is particularly useful when a node has multiple neighbors with different edge types or labels. In this case, the local node positional encoding can be learned based on these edge features:
$\textbf{p}_{i}=[f(\textbf{e}_{ij_{1}}),f(\textbf{e}_{ij_{2}}),...,f(\textbf{e}_{ij_{l}})],$ (6)
where $\textbf{e}_{ij}$ is edge feature between node $v_{i}$ and node $v_{j}$, $f$ is a learnable function that maps edge features to embeddings and $l$ is the number of neighbors considered.
To enhance the implementation of local node positional encodings in a broader context, a viable strategy is leveraging graph kernels or similarity functions to evaluate the structural similarity between two nodes within the graph [45, 37]. For instance, when a node exhibits three neighbors with unique subgraph patterns or motifs in their vicinity, its local node positional encoding can be computed as a vector of kernel values between the node and its neighboring nodes:
$\textbf{p}_{i}=[K(G_{i},G_{j_{1}}),K(G_{i},G_{j_{2}}),...,K(G_{i},G_{j_{l}})].$ (7)
In this equation, $G_{i}$ refers to the subgraph formed by node $v_{i}$ and its neighbors. The function $K$ is a graph kernel function that measures the similarity between two subgraphs. Mialon et al. [46] utilized this approach in their GraphiT model, which incorporates positive definite kernels on graphs as relative positional encodings for graph transformers.
Local node positional encodings enhance the self-attention mechanism of graph transformers by integrating structural and topological information of nodes. This encoding approach offers the advantage of preserving the sparsity and locality of graph structure, resulting in enhanced efficiency and interpretability. However, a limitation of this method is its restricted ability to capture long-range dependencies or global properties of graphs, which are essential for tasks such as graph matching or alignment.
Global Node Positional Encodings. The concept of global node positional encodings is inspired by the use of absolute positional encodings in NLP [47]. In NLP, as mentioned previously, every token receives a feature vector indicating its position within a sequence. Extending this idea to graph transformers, each node can be assigned a feature vector representing its position within the embedding space of the graph [48]. The objective of this encoding technique is to encapsulate the overall geometry and spectrum of graphs, thereby unveiling its intrinsic properties and characteristics.
One method for obtaining global node positional encodings is to leverage eigenvectors or eigenvalues of a matrix representation, such as adjacency matrix or Laplacian matrix [39]. For instance, if a node’s coordinates lie within the first k eigenvectors of graph Laplacian, its global node positional encoding can be represented by the coordinate vector:
$\textbf{p}_{i}=[u_{i1},u_{i2},...,u_{ik}],$ (8)
where $u_{ij}$ is $j$-th component of $i$-th eigenvector of the graph Laplacian matrix. One alternative approach to incorporating global node positional encodings is by utilizing diffusion or random walk techniques, such as personalized PageRank or heat kernel [49]. For example, if a node possesses a probability distribution over all other nodes in the graph, following a random walk [50], its global node positional encoding can be represented by this probability vector:
$\textbf{p}_{i}=[\pi_{i1},\pi_{i2},...,\pi_{iN}],$ (9)
where $\pi_{ij}$ is the probability of reaching node $v_{j}$ from node $v_{i}$ after performing a random walk on the graph.
A more prevalent approach to implementing global node positional encodings is to utilize graph embedding or dimensionality reduction techniques that map nodes to a lower-dimensional space while maintaining a sense of similarity or distance [51]. For instance, if a node possesses coordinates in a space derived from applying multi-dimensional scaling or graph neural networks to the graph, its global node positional encoding can be represented by that coordinate vector:
$\textbf{p}_{i}=[y_{i1},y_{i2},...,y_{ik}],$ (10)
where $y_{ij}$ is $j$-th component of $i$-th node embedding in a $k$-dimensional space, which can be obtained by minimizing the objective function that preserves graph structure:
$\min_{\textbf{Y}}\sum_{i,j=1}^{N}\textbf{w}_{ij}\|\textbf{y}_{i}-\textbf{y}_{j}\|^{2}.$ (11)
Here, $\textbf{w}_{ij}$ is a weight matrix that reflects the similarity or distance between node $v_{i}$ and node $v_{j}$, $Y$ is the matrix of node embeddings.
The primary aim of global node positional encodings is to improve the representation of node attributes in graph transformers by incorporating geometric and spectral information from graphs. This encoding method offers the advantage of capturing long-range dependencies and overall graph characteristics, benefiting tasks like graph matching and alignment. However, a drawback of this encoding approach is that it could undermine the sparsity and locality of graph structures, potentially impacting efficiency and interpretability.
III-A2 Edge Structural Bias
In the realm of graph transformers, edge structural bias is crucial for extracting and understanding complex information within graph structure [52]. Edge structural bias is versatile and can represent various aspects of graph structure, including node distances, edge types, edge directions and local sub-structures. Empirical evidence has shown that edge structural encodings can improve the effectiveness of graph transformers [53, 54, 55].
Local Edge Structural Encodings. Local edge structural encodings capture the local structure of a graph by encoding relative position or distance between two nodes [52]. These encodings borrow ideas from relative positional encodings used in NLP and CV, where they are used for modeling sequential or spatial order of tokens or pixels [56]. However, in the context of graphs, the concept of relative position or distance between nodes becomes ambiguous due to the presence of multiple connecting paths with varying lengths or weights. Consequently, the scientific community has proposed various methods to define and encode this information specifically for graph structures.
GraphiT [46] introduces local edge structural encodings to graph transformers. It promotes the use of positive definite kernels on graphs to measure node similarity considering the shortest path distance between them. The kernel function is defined as:
$k(u,v)=\exp(-\alpha d(u,v)),$ (12)
where $u$ and $v$ are two nodes in a graph, $d(u,v)$ is their shortest path distance and $\alpha$ is a hyperparameter that controls decay rate. The kernel function is then used to modify the self-attention score between two nodes as:
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+k(\textbf{Q},\textbf{K}))\textbf{V},$ (13)
where $k(\textbf{Q},\textbf{K})$ is the matrix of kernel values computed for each pair of nodes. EdgeBERT [57] proposes using edge features as additional input tokens for graph transformers. These edge features are obtained by applying a learnable function to the source and target node features of each edge. The resulting edge features are then concatenated with node features and fed into a standard transformer encoder.
More recently, the Edge-augmented Graph Transformer (EGT) [58] introduces residual edge channels as a mechanism to directly process and output both structural and node information. The residual edge channels are matrices that store edge information for each pair of nodes. They are initialized with either an adjacency matrix or the shortest path matrix and updated at each transformer layer by applying residual connections. These channels are then used to adjust the self-attention score between two nodes
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V},\textbf{R}_{e})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+\textbf{R}_{e})\textbf{V},$ (14)
where $\textbf{R}_{e}$ is the residual edge channel matrix.
Although local edge structural encodings can capture detailed structural information, they may have limitations in capturing overall structural information. This can lead to increased computational complexity or memory usage as pairwise information needs to be computed and stored. Additionally, the effectiveness of these encodings may vary depending on the selection and optimization of encoding or kernel function for different graphs and tasks.
Global Edge Structural Encodings. Global edge structural encodings aim to capture the overall structure of a graph. Unlike NLP and CV domains, the exact position of a node in graphs is not well-defined because there is no natural order or coordinate system [59]. Several approaches have been suggested to tackle this issue.
GPT-GNN [60] is an early work that utilizes graph pooling and unpooling operations to encode the hierarchical structure of a graph. It reduces graph size by grouping similar nodes and then restoring the original size by assigning cluster features to individual nodes. This approach results in a multiscale representation of graphs and has demonstrated enhanced performance on diverse tasks. Graphormer [11] uses spectral graph theory to encode global structure. It uses eigenvectors of normalized Laplacian matrix as global positional encodings for nodes. This method can capture global spectral features (e.g., connectivity, centrality and community structure). Park et al. [41] extended Graphormer by using singular value decomposition (SVD) to encode global structure. They utilized the left singular matrix of the adjacency matrix as global positional encodings for nodes. This approach can handle both symmetric and asymmetric matrices.
Global edge structural encodings excel at capturing coarse-grained structural information at the graph level, benefiting tasks that require global understanding. However, they may struggle with capturing fine-grained node-level information and can lose data or introduce noise during encoding. In addition, their effectiveness may depend on the choice of encoding technique and matrix representations.
III-A3 Message-passing Bias
Message-passing bias is a crucial inductive bias for graph transformers to facilitate learning from the local structure of graphs [61]. This bias enables graph transformers to exchange information between nodes and edges, thereby capturing dependencies and interactions among graph elements. Moreover, message-passing bias helps graph transformers in overcoming certain limitations of standard transformer architecture, such as the quadratic complexity of self-attention, the absence of positional information and challenges associated with handling sparse graphs [39]. Formally, message-passing bias can be expressed as follows:
$\textbf{h}_{v}^{(t+1)}=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$ (15)
Here, $\textbf{h}_{v}^{(t)}$ is the representation of node $v$ at layer $t$, $u$ is a neighboring node of node $v$, $\textbf{e}_{uv}$ is the feature of the edge between nodes $v$ and $u$, and $f$ is the function that aggregates information from neighbors and edges. Incorporation of message-passing bias in graph transformers can be achieved through three primary approaches: preprocessing, interleaving and post-processing (as illustrated in Figure 4 ). These approaches vary in how they combine message-passing operations with self-attention layers within transformer architecture.
Figure 4: Incorporation of message-passing bias in transformer model.
Preprocessing. Preprocessing involves applying message-passing operations to node features before feeding them to self-attention layers [36]. This technique aims to augment node features with local structural information, making them more compatible with global self-attention. Preprocessing can maintain the integrity of the original transformer architecture without making any modifications and utilize message-passing modules from existing GNNs. It can be mathematically defined as follows:
$\textbf{h}_{v}^{(t)}$ $=f(\textbf{h}_{v}^{(t-1)},\{\textbf{h}_{u}^{(t-1)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$ (16)
$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}).$
Here, $\textbf{h}_{v}^{(0)}=\textbf{x}_{v}$, $\textbf{x}_{v}$ is input feature of node $v$, SelfAttention is a function that performs self-attention over all nodes.
The limitation of preprocessing is that it applies message-passing only once before self-attention layers. This approach may not fully capture intricate interactions between nodes and edges on various scales. Additionally, the preprocessing step may introduce redundancy and inconsistency between the message-passing module and the self-attention layer as they both serve similar functions of aggregating information from neighboring elements.
Interleaving. Interleaving refers to the technique employed in graph transformer architecture that involves alternating message-passing operations and self-attention layers [62, 16, 63]. The objective of this technique is to achieve a balance between local and global information processing, thereby enabling multi-hop reasoning over graphs. By integrating message-passing modules into core components of graph transformers, interleaving enhances their expressive power and flexibility. It can be mathematically defined as follows:
$\textbf{h}_{v}^{(t+1)}$ $=\theta+\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$ (17)
$\theta$ $=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
One drawback of the interleaving technique is its impact on the complexity and computational requirements of graph transformers. This is due to the need for additional parameters and operations compared to pre-processing or post-processing. Furthermore, interleaving can potentially lead to interference and conflict between the message-passing module and the self-attention layer as they each update node representations in distinct manners.
Post-processing. Post-processing refers to the technique of applying message-passing operations to node representations obtained from self-attention layers [64, 58]. The purpose of this approach is to refine and adjust node representations based on underlying graph structure, thereby enhancing their interpretability and robustness. By doing so, this method aims to improve the quality and utility of node representations for downstream tasks and applications. It can be mathematically defined as follows:
$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$ (18)
$\textbf{h}_{v}^{(T+1)}$ $=f(\textbf{h}_{v}^{(T)},\{\textbf{h}_{u}^{(T)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
Here, $T$ is the final layer of the graph transformer.
The drawback of post-processing is its limited application of message-passing after self-attention layers, potentially failing to capture intricate semantics and dynamics of graph data. Additionally, post-processing runs the risk of introducing noise and distortion to node representations as it has the potential to overwrite or override information acquired by self-attention layers.
III-A4 Attention Bias
Attention bias enables graph transformers to effectively incorporate graph structure information into the attention mechanism without message-passing or positional encodings [39, 65]. Attention bias modifies attention scores between nodes based on their relative positions or distances in the graph. It can be categorized as either local or global depending on whether it focuses on the local neighborhood or global topology of the graph.
Local Attention Bias. Local attention bias limits the attention to a local neighborhood surrounding each node. This concept is analogous to the message-passing mechanism observed in GNNs [66]. It can be mathematically defined as follows:
[…]
where $\alpha_{ij}$ is the attention score between node $v_{i}$ and node $v_{j}$, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are their node features, $g$ is a function that computes the similarity between two nodes, such as dot-product and linear transformation. $b_{ij}$ is a local attention bias term that modifies attention score based on the distance between node $v_{i}$ and node $v_{j}$. The local attention bias term can be either a binary mask that only allows attention within a certain hop distance [67, 42, 36] or a decay function that decreases attention score with increasing distance [68, 64].
Global Attention Bias. Global attention bias integrates global topology information into the attention mechanism independent of message-passing and positional encodings [69]. It can be mathematically defined as follows:
$\alpha_{ij}=\frac{\exp(g(\mathbf{x}_{i},\mathbf{x}_{j})+c(\mathbf{A},\mathbf{D},\mathbf{L},\mathbf{P})_{ij})}{\sum_{k=1}^{N}\exp(g(\mathbf{x}_{i},\mathbf{x}_{k})+c(\mathbf{A},\mathbf{D},\mathbf{L},\mathbf{P})_{ik})}.$ (20)
Here, $c$ is the function that computes the global attention bias term, modifying attention score based on some graph-specific matrices or vectors, such as adjacency matrix $\mathbf{A}$, degree matrix $\mathbf{D}$, Laplacian matrix $\mathbf{L}$, and PageRank vector $\mathbf{P}$. The global attention bias term can be additive or multiplicative to the similarity function [46, 70]. Generally, global attention bias can enhance the global structure awareness and expressive power of graph transformers [70].
III-B Graph Attention Mechanisms
Graph attention mechanisms play an important role in the construction of graph transformers [71]. By dynamically assigning varying weights to nodes and edges in the graph, these mechanisms enable transformers to prioritize and emphasize the most relevant and important elements for a given task [72]. Specifically, a graph attention mechanism is a function that maps each node $v_{i}\in V$ to a vector $\textbf{h}_{i}\in\mathbb{R}^{d_{k}}$:
$\textbf{h}_{i}=f_{n}(\textbf{x}_{i},\{\textbf{x}_{j}\}_{v_{j}\in\mathcal{N}(v_{i})},\{A_{ij}\}_{v_{j}\in\mathcal{N}(v_{i})}),$ (21)
where $f_{n}$ is a nonlinear transformation, $\textbf{x}_{i}$ is the input feature vector of node $v_{i}$, $\textbf{x}_{j}$ is the input feature vector of node $v_{j}$. The function $f_{n}$ can be decomposed into two parts: an attention function and an aggregation function.
The attention function computes a scalar weight for each neighbor of node $v_{i}$, denoted by $\alpha_{ij}$, which reflects the importance or relevance of node $v_{j}$ for node $v_{i}$:
$\alpha_{ij}=\text{softmax}_{i}(\text{LeakyReLU}(\textbf{W}_{a}[\textbf{x}_{i}||\textbf{x}_{j}])),$ (22)
where $\textbf{W}_{a}\in\mathbb{R}^{1\times 2d_{n}}$ is a learnable weight matrix, ||||| | is the concatenation operation, and $\text{softmax}_{i}$normalizes weights over all neighbors of node $v_{i}$. The aggregation function combines the weighted features of the neighbors to obtain the output representation of node $v_{i}$:
$\textbf{h}_{i}=\textbf{W}_{h}\left(\textbf{x}_{i}+\sum_{v_{j}\in\mathcal{N}(v_{i})}\alpha_{ij}\textbf{x}_{j}\right),$ (23)
where $\textbf{W}_{h}$ is another learnable weight matrix. Graph attention mechanisms have the potential to extend their application beyond nodes to edges. This can be achieved by substituting nodes with edges and utilizing edge features instead of node features [73]. Furthermore, stacking of graph attention mechanisms in multiple layers allows for the utilization of output representations from one layer as input features for the subsequent layer [74].
III-B1 Global Attention Mechanisms
Global attention mechanisms can determine how each node calculates its attention weights across all other nodes in the graph [75]. Global attention mechanisms can be broadly categorized into two types: quadratic attention mechanisms and linear attention mechanisms.
Quadratic Attention Mechanisms. Quadratic attention mechanisms are derived from the conventional self-attention formula. This formula calculates attention weights by applying a softmax function to scale the dot product between the query and key vectors of each node:
$\alpha_{ij}=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})}.$ (24)
The computational complexity of this method is $O(N^{2})$.
One of the pioneering studies that has introduced quadratic attention mechanisms for graph transformers is the work by Velickovic et al.[42]. The authors proposed the use of multiple attention heads to capture different types of relations between nodes. Choromanski et al.[76] introduced the Graph Kernel Attention Transformer (GKAT), an approach that integrates graph kernels, structural priors, and efficient transformer architectures. GKAT leverages graph kernels as positional encodings to capture the structural characteristics of the graph while employing low-rank decomposition techniques to minimize the memory requirements of quadratic attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{j})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{n})},$ (25)
$\mathbf{z}_{i}$ $=\sum_{j=1}^{N}\alpha_{ij}(\mathbf{U}\mathbf{V}^{\top})_{ij}$
where $\textbf{z}_{i}$ is the output vector of node $v_{i}$, $\mathbf{p}_{i}$ is the positional encoding vector of node $v_{i}$, which is computed by applying a graph kernel function to node features. $\mathbf{U}$ and $\mathbf{V}$ are low-rank matrices that approximate value matrix in the GKAT model. GKAT also introduced a novel kernel-based masking scheme to control the sparsity of the attention matrix. Yun et al. [36] introduced Graph Transformer Networks (GTN), an approach that uses multiple layers of quadratic attention to acquire hierarchical representations of graphs. GTN further enhances attention computation by incorporating edge features through the use of bilinear transformations instead of concatenation. This innovative technique demonstrates the potential for improved graph representation learning in the field of graph analysis.
Quadratic attention mechanisms possess a significant advantage in capturing extensive dependencies and global information within graphs, thereby offering potential benefits for various graph learning tasks. Nonetheless, these mechanisms have some limitations, including their computationally expensive nature, high memory consumption, and susceptibility to noise and outliers. Additionally, quadratic attention mechanisms may not effectively preserve and leverage local information within graphs which holds significance in certain domains and tasks.
Linear Attention Mechanisms. Linear attention mechanisms employ different approximations and hashing techniques to decrease the computational complexity of self-attention from $O(N^{2})$ to $O(N)$ [77]. These mechanisms can be categorized into two subtypes: kernel-based linear attention mechanisms and locality-sensitive linear attention mechanisms.
Kernel-based linear attention mechanisms leverage the concept of employing kernel functions to map query and key vectors into a feature space that allows efficient computation of their inner products [78, 79]. In the work by Katharopoulos et al. [80], Linear Transformer Networks (LTN) were introduced. LTN utilizes random Fourier features as kernel functions to approximate self-attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{j})}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{n})}{\sqrt{d_{k}}})},$ (26)
$\phi(\mathbf{x})$ $=\sqrt{\frac{2}{m}}[\cos(\omega_{1}^{\top}\mathbf{x}+b_{1}),...,\cos(\omega_{m}^{\top}\mathbf{x}+b_{m})]^{\top}.$
Here, $\phi$ is a random Fourier feature function that maps input vector into a $m$-dimensional feature space. $\omega_{i}$ and $b_{i}$ are randomly sampled from a Gaussian distribution and a uniform distribution, respectively.
Locality-sensitive linear attention mechanisms are based on the concept of employing hashing functions to partition the query and key vectors into discrete buckets that facilitate local computation of their inner products [81, 82, 83]. Kitaev et al. [84] introduced the Reformer model which uses locality-sensitive hashing (LSH) as a hashing function to approximate self-attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n\in B_{i}}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})},$ (27)
$B_{i}$ =$=\{j|h(\mathbf{q}_{i})=h(\mathbf{k}_{j})\},$
where $h$ is an LSH function that maps the input vector into a bucket index and $B_{i}$ is a set of nodes that share the same bucket index with node $v_{i}$. RGN also incorporated edge features into the attention computation using a shared query-key projection matrix.
Linear attention mechanisms offer a significant advantage in terms of reducing computational cost and memory usage in self-attention, thereby enabling graph transformers to efficiently process large-scale graphs. However, they do come with certain limitations, including approximation error, hashing collision and loss of global information. Additionally, linear attention mechanisms may not be capable of capturing intricate or nonlinear relationships between nodes.
III-B2 Local Attention Mechanisms
Local attention mechanisms determine how each node computes its attention weights over a subset of nodes in the graph [85]. Local attention mechanisms can be broadly categorized into two types: message-passing attention mechanisms and spectral attention mechanisms.
Message-passing Attention Mechanisms. Message-passing attention mechanisms are built on the foundational framework of message-passing neural networks (MPNNs) [86], which iteratively aggregate messages from neighboring nodes to compute node representations. To enhance MPNNs, message-passing attention mechanisms employ self-attention to compute messages or aggregation weights [61, 87, 88, 39]. The computational complexity of this method is $O(E)$, where $E$ represents the number of edges in the graph.
One of the pioneering works in introducing message-passing attention mechanisms for graph transformers is GraphSAGE [89]. GraphSAGE proposes the use of mean pooling as an aggregation function and self-attention as a combination function. Additionally, GraphSAGE incorporates node sampling techniques to efficiently handle large-scale graphs. The function can be mathematically defined as:
$\mathbf{h}_{i}^{(l+1)}$ $=\text{ReLU}(\mathbf{W}^{(l)}[\mathbf{h}_{i}^{(l)}\|\text{MEAN}(\{\mathbf{h}_{j}^{(l)},\forall v_{j}\in\mathcal{N}(v_{i})\})])$ (28)
$\mathbf{z}_{i}$ $=\sum_{l=0}^{L}\alpha_{il}\mathbf{h}_{i}^{(l)}.$
[…]
Spectral Attention Mechanisms. Spectral attention mechanisms are founded on the concept of transforming node features into a spectral domain, where the graph structure is encoded by eigenvalues and eigenvectors of the graph Laplacian matrix [91]. Spectral attention mechanisms employ self-attention to calculate spectral coefficients and spectral filters [92]. The computational complexity of this approach is $O(N)$.
Wang et al. [22] proposed graph isomorphism networks (GIN) as a spectral attention mechanism for graph transformers, employing an expressive aggregation function of sum pooling and self-attention. GIN also integrates a unique graph readout scheme via set functions to encapsulate global graph characteristics. The equations of GIN are as follows:
$\mathbf{h}_{i}^{(l+1)}$ $=\text{MLP}^{(l)}((1+\epsilon^{(l)})\mathbf{h}_{i}^{(l)}+\sum_{j\in\mathcal{N}(v_{i})}\mathbf{h}_{j}^{(l)}),$ (29)
$\mathbf{z}_{i}$ $=\sum_{l=0}^{L}\alpha_{il}\mathbf{h}_{i}^{(l)},$
$\mathbf{z}_{G}$ $=\text{READOUT}(\{\mathbf{z}_{i},\forall v_{i}\in V\}).$
Here, $\text{MLP}^{(l)}$ is a multi-layer perceptron, $\epsilon^{(l)}$ is a learnable parameter, READOUT indicates a set function that aggregates node output vectors into a graph output vector. $\mathbf{z}_{G}$ is the output vector of graph $G$. In addition, Nguyen et al. [93] introduced UGformer, a self-attention-based method for the spectral coefficient computation of each node using eigenvalues and eigenvectors of the Laplacian graph matrix. UGformer further integrates edge features into spectral computation via bilinear transformations.
Spectral attention mechanisms possess the ability to incorporate the structural information of a graph into the spectral domain, thereby offering potential benefits for certain tasks or domains. However, they are also accompanied by certain limitations, including high computational cost, memory consumption and sensitivity to graph size and density.
IV Taxonomy of Graph Transformers
The past few years have witnessed a surge of interest in graph transformers. This section dives into four key categories dominating the current literature: shallow, deep, scalable, and pre-trained graph transformers. By analyzing representative models within each category, we aim to establish valuable guidelines for designing effective graph transformers.
IV-A Shallow Graph Transformers
Shallow graph transformers represent a class of GNNs that leverage the power of self-attention to acquire node representations from data structured in graphs. Inspired by transformer models, which effectively capture long-range dependencies in sequential data through self-attention, shallow graph transformers extend this concept to graph data by computing self-attention weights based on both node features and graph topology [94]. The primary objective of shallow graph transformers is to achieve exceptional performance while minimizing computational complexity and memory usage.
Shallow graph transformers can be seen as a generalization of graph attention networks (GAT) [42]. GAT use a multi-head attention mechanism to calculate node embeddings. However, GAT has some limitations, such as the inability to model edge features and lack of diversity among attention heads [95]. Several GAT extensions have been proposed in the literature to address these issues. For example, GTN by Yun et al. [36] introduces edge-wise self-attention to incorporate edge information into node embeddings. Ahmad et al.[63] proposed the graph attention transformer encoder (GATE), which applies a masked self-attention mechanism to learn different attention patterns for different nodes. GATE also uses a position-wise feed-forward network and dropout to enhance model capacity and generalization. The summary of shallow graph transformer methods is given in Table II.
Shallow graph transformers are efficient and adaptable capable of handling various graph learning tasks and different types of graphs, but their lack of depth and recurrence may limit their ability to capture complex dependencies. Their performance can also be influenced by the choice of mask matrix and the number of attention heads, indicating a need for further research on their optimal design and regularization.
TABLE II: Shallow graph transformers
Model Graph inductive bias Graph attention mechanisms Application Dataset Code availability
Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
Graph Transformer [16] Yes (local) Yes (local) Yes (pre-processing) No Yes (message passing) Graph representation learning ZINC, PATTERN, CLUSTER Yes
GAT [42] No No No Yes (linear) Yes (message passing) Node classification of graph-structured data Cora, Citeseer, Pubmed, PPI Yes
GATv2 [95] No No No Yes (linear) Yes (message passing) Graph representation learning OGB, QM9, VARMISUSE Yes
GTN [36] No No No No Yes (message passing) Node classification on heterogeneous graphs DBLP, ACM, IMDB Yes
GATE [63] No No No Yes (linear) No Cross-lingual relation and event extraction ACE 2005 corpus Yes
PiT [96] No No No Yes (quadratic) No Image classification, object detection ImageNet, COCO, ImageNet-A Yes
WDPruning [97] No No No Yes (quadratic) No Image classification and image retrieval CIFAR-10 and ILSVRC-12 No
SVT [98] No No No Yes (quadratic) No Image classification Synthetic data and CIFAR-10 No
Gubelmann et al. [99] No No No Yes No Natural language inference (NLI) SNLI and MNLI No
TADDY [100] Yes (global and local) No No Yes (linear) No Anomaly detection in dynamic graphs UCI Messages, Enron, DBLP, Facebook, Twitter and Reddit. No
Tree Transformer [38] Yes (local) No No Yes (quadratic) No Program translation and semantic parsing For2Lam, CoffeeScript-JavaScript, JOBS, GEO, ATIS Yes
GAT-POS  [43] Yes (local) No No Yes (linear) Yes (message passing) Node classification Chameleon, Squirrel, Actor, Cora, Citeseer, Pubmed No
DGT [45] Yes No No Yes (linear) No Node classification on graphs Actor, Squirrel, Chameleon, Cora, Citeseer, twitch-gamers, ogbn-arxiv, Reddit No
EDTER [54] No No No Yes (linear) Yes (message passing) Edge detection BSDS500, NYUDv2, Multicue Yes
BGT-Net [55] No No No Yes No Scene Graph Generation for Images Visual Genome (VG), OpenImages (OI), Visual Relationship Detection (VRD) No
PMPNet [61] No No Yes (pre-processing) No Yes (message passing) 3D object detection nuScenes Yes
GAMLP [74] No No No Yes (linear) Yes (spectral) Node classification 14 real-world graph datasets Yes
SA-GAT [101] No No No Yes (linear) Yes (spectral) Graph classification MUTAG, PTC, NCI1, NCI109, PROTEINS, ENZYMES, DD and COLLAB No
GKAT [76] No No No Yes (linear) No Graph classification and graph pattern detection Bioinformatics and social networks datasets, ImageNet dataset No
EcoFormer [78] No No No Yes (linear) No Image classification and long-range tasks CIFAR-100, ImageNet-1K, Long Range Arena Yes
Reformer [84] No No No Yes (linear) Yes (message passing) NLP and generative tasks Synthetic task, enwik8, imagenet64, WMT 2014 English-German Yes
LISA [85] No No No Yes (linear) No Sequential recommendation Alibaba, Amazon Video Games, MovieLens 1M, MovieLens 25M Yes
IV-B Deep Graph Transformers
Deep graph transformers consist of multiple self-attention layers stacked on top of each other, with optional skip connections, residual connections or dense connections between layers [102]. They are designed to achieve higher performance with increased model depth and complexity [103]. Deep graph transformers extend shallow graph transformers by applying self-attention layers to node features and graph topology hierarchically.
However, deep graph transformers also face several challenges that need to be addressed. One challenge is the difficulty of training deeper models, which can be mitigated by employing techniques, such as PairNorm introduced in DeeperGCN [104]. Another challenge is the over-smoothing problem, which can be addressed by using a gated residual connection and a generalized convolution operator as proposed in DeeperGCN. Additionally, the disappearance of global attention capacity and the lack of diversity among attention heads are challenges that can be tackled by approaches like DeepGraph [94]. DeepGraph incorporates substructure tokens and local attention to improve the focus and diversity of global attention.
Deep graph transformers, while complex, can achieve top-tier results on various graph learning tasks and adapt to different types of graphs and domains. However, their high computational cost, difficulty in optimization, and sensitivity to hyperparameters necessitate further research for optimal design and training. Table III summarizes the methods of deep graph transformers.
TABLE III: Deep graph transformers
Model Graph inductive bias Graph attention mechanisms Application Datasets Code availability
Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
HG-Transformer [102] No No No Yes (linear) No Multi-label text classification RCV1, RCV1-2K, AmazonCat-13K No
Unfolded Transformer [103] No No No Yes (quadratic) No Graph neural networks and NLP IMDB and SST2 Yes
DeepGraph [94] Yes (relative) No No Yes (linear) Yes (message passing) Graph representation learning PCQM4M-LSC, ZINC, CLUSTER, PATTERN Yes
GraphWriter [105] No Yes No Yes (linear) No Text generation from knowledge graphs AGENDA dataset Yes
GraphLoG [106] No No Yes (pre-processing) No Yes (message passing) Graph representation learning ZINC15, MoleculeNet, Protein ego-networks Yes
DIET [40] Yes No No Yes (quadratic) No cross-lingual generalization, machine translation GLUE, XTREME, WMT 2018 No
GRPE [41] Yes (local and global) Yes (local and global) No Yes (linear) No Graph representation learning ZINC, MolPCBA, MolHIV, PATTERN, CLUSTER, PCQM4M, PCQM4Mv2 No
Graphormer [11] No Yes No Yes (quadratic) No Graph representation learning PCQM4M-LSC, OGB, ZINC Yes
Relation-aware Self-Attention [26] No Yes (local) No Yes (quadratic) No Machine translation WMT 2014 English-German and English-French Yes
GraphiT [46] Yes (local and global) No No Yes (quadratic) Yes (message passing) Graph representation learning for classification and regression tasks MUTAG, PROTEINS, PTC, NCI1, ZINC Yes
Tree Transformer [48] Yes (global and local) No No Yes (quadratic) Yes (message passing) Code representation learning Python150k and JavaScript150k Yes
MagLapNet [49] Yes (local and global) No No Yes (quadratic) Yes (spectral) Graph learning tasks such as correctness testing of sorting networks and source code understanding Synthetic graphs, OGB Code2 Yes
SAT [52] Yes (local and global) No No Yes (quadratic) Yes (message passing) Graph representation learning for graph and node property prediction ZINC, CLUSTER, PATTERN, OGBG-PPA, OGBG-CODE2 Yes
Edge Transformer [107] Yes Yes No Yes (quadratic) No Improve systematic generalization in natural language understanding and relational reasoning tasks CLUTRR, CFQ and COGS Yes
Cai et al. [62] Yes Yes No Yes (quadratic) No Graph-to-sequence learning LDC2015E86, LDC2017T10, WMT16 English-German, WMT16 English-Czech Yes
SAN [66] Yes (local and global) No No Yes (quadratic) Yes (spectral) Graph representation learning ZINC, PATTERN, CLUSTER, MolHIV, MolPCBA Yes
KMCA [79] No No No Yes (quadratic) Yes (message passing) Time series analysis Synthetic, FordA, uWave, IMDB, ECG No
ABT-MPNN [87] No No Yes (pre-processing) Yes (quadratic and linear) No Molecular property prediction and biological activity prediction Johnson et al. [108], Tox21, ClinTox, ToxCast, HIV, QM8, ESOL, FreeSolv, Lipophilicity and QM8. Yes
UGformer [93] No No No Yes (quadratic) No Graph classification and text classification COLLAB, IMDB-B, IMDB-M, DD, PROTEINS, MUTAG, PTC, MR, R8, R52, Ohsumed Yes
DTI-GTN [109] No Yes Yes (interleaving) Yes (quadratic) No Drug-target interaction prediction Peng et al. [110] dataset No
Dual-GCN+Transformer [111] No No No Yes (linear) No Image Captioning MS COCO and Visual Genome Yes
SGTransformer [112] Yes Yes No Yes (quadratic) Yes (message passing) Scene graph to image generation Visual Genome, COCO-Stuff and CLEVR-Dialog Yes,
TPT [113] Yes No No Yes No Video scene graph generation VidHOI and Action Genome No
IV-C Scalable Graph Transformers
Scalable graph transformers are a category of graph transformers that tackle the challenges of scalability and efficiency when applying self-attention to large-scale graphs [114, 39, 53, 115]. These transformers are specifically designed to reduce computational cost and memory usage while maintaining or improving performance. To achieve this, various techniques are employed to reduce the complexity of self-attention, such as sparse attention, local attention, and low-rank approximation [12, 115]. Scalable graph transformers can be regarded as an enhancement of deep graph transformers addressing challenges, such as over-smoothing and limited capacity of global attention.
Several scalable graph transformer models have been proposed to enhance the scalability and efficiency of graph transformers. For instance, Rampášek et al. [39] introduced GPS, use low-rank matrix approximations to reduce computational complexity, and achieve state-of-the-art results on diverse benchmarks. GPS decouples local real-edge aggregation from a fully-connected transformer and incorporates different positional and structural encodings to capture graph topology. It also offers a modular framework that supports multiple encoding types and mechanisms for local and global attention. Cong et al. [116] developed DyFormer, a dynamic graph transformer that utilizes substructure tokens and local attention to enhance the focus and diversity of global attention. DyFormer employs a temporal union graph structure and a subgraph-based node sampling strategy for efficient and scalable training.
Scalable graph transformers are an innovative and efficient category of graph transformers that excel in handling large-scale graphs while minimizing computational cost and memory usage. However, scalable graph transformers face certain limitations, including the trade-off between scalability and expressiveness, the challenge of selecting optimal hyperparameters and encodings, and the absence of theoretical analysis regarding their convergence and stability. Consequently, further investigation is required to explore optimal designs and evaluations of scalable graph transformers for various applications. For a comprehensive overview of scalable graph transformer methods, please refer to Table IV.
TABLE IV: Scalable graph transformers
Model Graph inductive bias Graph attention mechanisms Application Datasets Code availability
Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
GPS [39] Yes (local and global) Yes (local and relative) Yes (interleaving) Yes (quadratic and linear) Yes (message passing) Graph representation learning ZINC, PCQM4Mv2, CIFAR10, MalNet-Tiny, OGB benchmarks Yes
NodeFormer [53] No No Yes (relational bias) Yes (linear) No Node classification Citation networks, OGB-Proteins, Amazon2M, Mini-ImageNet, 20News-Groups Yes
DIFFORMER [115] No No No Yes (linear) Yes (spectral) Node classification, image/text classification, spatial-temporal dynamics prediction Cora, Citeseer, Pubmed, ogbn-Proteins, Pokec, CIFAR, STL, 20News Yes
DyFormer [116] Yes (global) Yes (global) Yes (pre-processing) Yes (quadratic) No Dynamic graph representation learning RDS, UCI, Yelp, ML-10M, SNAP-Wikipedia, SNAP-Reddit No
Graph ViT/MLP-Mixer [117] Yes No No No Yes (message passing) Graph representation learning ZINC, MNIST, CIFAR10, MolTOX21, MolHIV, Peptide-func, Peptide-struct, CSL, EXP, SR25, LRGB, TreeNeighbourMatch Yes
DiGress [118] No Yes No Yes Yes Graph generation QM9, MOSES, GuacaMol, SBM, planar graphs Yes
EXPHORMER [119] Yes (local and global) Yes (local and global) Yes (interleaving) Yes (linear) Yes (spectral) Graph learning and representation tasks CIFAR10, MalNet-Tiny, MNIST, CLUSTER, PATTERN, PascalVOC-SP, COCO-SP, PCQM-Contact, ogbn-arxiv, Yes
EGT [120] Yes (global) Yes (global) No Yes (quadratic) No Node and edge classification, graph classification and regression, transfer learning. PATTERN, CLUSTER, TSP, MNIST, CIFAR10, ZINC, PCQM4M, PCQM4Mv2, MolPCBA and MolHIV. Yes
GRIT [121] Yes No No Yes (quadratic) No Graph regression and classification ZINC, MNIST, CIFAR10, PATTERN, CLUSTER, Peptides-func, Peptides-struct, ZINC-full, PCQM4Mv2 Yes
HSGT [122] No Yes (local, based on shortest path distance) No Yes (quadratic) Yes (message passing, based on neighborhood sampling) Node classification Cora, CiteSeer, PubMed, Amazon-Photo, ogbn-arxiv, ogbn-proteins, ogbn-products, Reddit, Flickr, Yelp No
Higher-Order Transformers [123] No No No Yes (quadratic and linear) Yes (message passing) Graph and hypergraph modeling Synthetic chains, PCQM4M-LSC, Jets, Delaunay, Hyperedge prediction Yes
EGAT [73] No No Yes (interleaving) Yes (linear) No Graph representation learning and molecular property prediction AMLSim, Cora, Citeseer, PubMed No
L-CAT [90] No No No Yes (linear) No Node classification CSBM, Amazon, Facebook, GitHub, TwitchEN, Coauthor Physics, OGB (arxiv, products, mag, proteins) Yes
Point Transformer [88] Yes (relative) No No No Yes (message passing) 3D point cloud processing S3DIS, ModelNet40, ShapeNetPart Yes
Specformer [92] No No No Yes (linear) Yes (spectral) Graph representation learning Synthetic, node-level and graph-level datasets Yes
HierGAT [124] Yes (local) No No Yes (linear) No Entity Resolution Magellan and WDC product matching datasets Yes
Geo-ER [125] No No No Yes (linear) Yes (message passing) Geospatial Entity Resolution OSM, FSQ, Yelp Yes
GTA [126] No No No Yes (linear) Yes (message passing) Multivariate time series anomaly detection in IoT SWaT, WADI, SMAP, MSL Yes
GraTransDRP [127] Yes No No Yes (linear) No Drug response prediction GDSC Yes
LGT [128] Yes (local and global) No No Yes (linear) Yes (message passing) Molecular properties prediction ZINC, QM9, ESOL, FreeSolv No,
GTAGC [129] Yes (global) No No Yes (linear) Yes (spectral) Graph clustering Citeseer, Cora, Pubmed No
ANS-GT [130] Yes (local and global) No No Yes (linear) Yes (message passing) Node classification Cora, Citeseer, Pubmed, Chameleon, Actor, Squirrel, Cornell, Texas, Wisconsin No
RelTR [131] No No No Yes (linear) No Scene graph generation Visual Genome, Open Images V6, VRD Yes
BSTG-Trans [132] Yes (local and global) Yes (local and global) No Yes (linear) Yes (message Passing) Long-term Pose Forecasting Human3.6M, HumanEva-I, Human360K Yes
GraphSum [133] Yes (local and global) No No Yes (linear) Yes (message passing) Abstractive MDS WikiSum and MultiNews Yes
KG-Transformer [134] No No Yes (pre-processing) Yes (linear) No Text generation from knowledge graphs UMLS, PubMed, 2010 i2b2/VA, JNLPBA, BC5CDR, GAD, EU-ADR No
IV-D Pre-trained Graph Transformers
Pre-trained graph transformers utilize large-scale unlabeled graphs to acquire transferable node embeddings [135]. These embeddings can be fine-tuned for downstream tasks with scarce labeled data that address the challenges of data scarcity and domain adaptation in graph learning tasks [136, 137]. These transformers are similar to pre-trained large language models (LLMs) and are trained on graph datasets using self-supervised learning objectives, such as masked node prediction [138], edge reconstruction [139], and graph contrastive learning [140]. These objectives aim to encapsulate the inherent properties of graph data independently of external labels or supervision [141]. The pre-trained model can be fine-tuned on a specific downstream task with a smaller or domain-specific graph dataset by incorporating a task-specific layer or loss function and optimizing it on labeled data [142]. This allows the pre-trained model to transfer the knowledge acquired from the large-scale graph dataset to the subsequent task, giving better performance compared to the training from scratch [142].
Pre-trained graph transformers face some challenges, such as the selection of appropriate pre-training tasks, domain knowledge incorporation, heterogeneous information integration, and pre-training quality evaluation [143]. To address these issues, KPGT [144] and KGTransformer [145] have been proposed. KPGT leverages additional domain knowledge for pre-training, while KGTransformer serves as a uniform Knowledge Representation and Fusion (KRF) module in diverse tasks. Despite their power and flexibility, pre-trained graph transformers encounter issues related to graph data heterogeneity and sparsity, domain adaptation, model generalization and performance interpretation. A summary of pre-trained graph transformer methods is provided in Table V.
TABLE V: Pre-trained graph transformers
Model Graph inductive bias Graph attention mechanisms Application Datasets Code availability
Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
GRAPHIX-T5 [136] No Yes Yes (pre-processing) Yes (linear) No Text-to-SQL SPIDER, SYN, DK, REALISTIC, SPIDER-SSP Yes
GROVER [137] No No Yes (interleaving) Yes (linear) Yes (message passing) Molecular representation learning MoleculeNet No
G-BERT [135] No No No Yes (linear) No Medication recommendation MIMIC-III Yes
ChemBERTa [138] No No No Yes (quadratic) No Molecular property prediction PubChem 77M, MoleculeNet Yes
HEAT [139] Yes (local) No No Yes (linear) Yes (message passing) Structured reconstruction SpaceNet Challenge and Structured3D Yes
FormNetV2 [140] No No No Yes (linear) Yes (message passing) Form document information extraction FUNSD, CORD, SROIE No
MPG [143] No No No Yes (linear) Yes (spectral) Drug discovery MoleculeNet, DDI, DTI Yes
LiGhT [144] No Yes (path encoding and distance encoding) No Yes (linear) No Molecular property prediction ChEMBL29, BACE, BBBP, ClinTox, SIDER, Estrogen, MetStab, Tox21, ToxCast, FreeSolv, Lipophilicity Yes
LightGT [146] Yes (global) No No Yes (linear) No Multimedia recommendation Movielens, Tiktok, Kwai Yes
KGTransformer [145] No No No Yes (linear) Yes (message passing) Knowledge graph transfer WFC, WN18RR, AwA-KG, CommonsenQA Yes
CoVGT [147] No No No Yes (linear) Yes (message-passing) Video Question Answering NExT-QA, TGIF-QA, TGIF-QA-R, STAR-QA, Causal-VidQA, MSRVTT-QA Yes
G-Adapter [148] No Yes (local) No Yes (quadratic) No Molecular graph tasks MolHIV, MolPCBA, FreeSolv, ESOL, BBBP, Estrogen-$\alpha$, Estrogen-$\beta$, MetStablow, MetStabhigh Yes
kgTransformer [149] No No Yes (interleaving) Yes (quadratic) No Knowledge graph reasoning for complex logical queries FB15k-237, NELL995 Yes
Li et al. [149] Yes (local and global) No No Yes (linear) Yes (message passing) Movement synchrony estimation in autism therapy interventions PT13 and Human3.6MS No
Pellegrini et al. [150] No Yes No Yes (linear) No Disease prediction using graph convolutional networks TADPOLE and MIMIC-III Yes
Video Graph Transformer (VGT) [151] No No No Yes (linear) Yes (message passing) Video Question Answering (VideoQA) NExT-QA, TGIF-QA, MSRVTT-QA Yes
Tan et al. [152] No No No Yes (linear) No Few-shot node classification CoraFull, ogbn-arxiv, CiteSeer, Cora No
GPT-GNN [60] No Yes Yes No Yes Graph representation learning and graph mining OAG, Amazon, Reddit, OAG (citation) Yes
IV-E Design Guide for Effective Graph Transformers
Developing effective graph transformers requires meticulous attention to detail and careful consideration. This guide provides general principles and tips for designing graph transformers for various scenarios and tasks.
  • Choose the appropriate type of graph transformers based on the nature and complexity of your graph data and tasks. For simple and small graph data, a shallow graph transformer with a few layers may suffice. For complex and large graph data, a deep graph transformer with many layers can learn more expressive representations. For dynamic or streaming graph data, a scalable graph transformer is more efficient. Pre-trained graph transformers are more suitable for sparse or noisy graph data.
  • Design suitable structural and positional encodings for your graph data. These encodings capture the structure of graphs and are added to input node or edge features before feeding them to transformer layers. The choice of encodings depends on the characteristics of the graph data, such as directionality, weight, and homogeneity. The careful design of these encodings ensures their informativeness.
  • Optimize the self-attention mechanism for your graph data. Self-attention mechanisms compute attention scores among all pairs of nodes or edges in the graph, capturing long-range dependencies and interactions. However, it introduces challenges like computational complexity, memory consumption, overfitting, over-smoothing, and over-squashing. Techniques like sampling, sparsification, partitioning, hashing, masking, regularization, and normalization can be employed to address these challenges and improve the quality and efficiency of the self-attention mechanism.
  • Utilize pre-training techniques to enhance the performance of graph transformers. Pre-training techniques leverage pre-trained models or data from other domains or tasks transferring knowledge or parameters to a specific graph learning task. Methods like fine-tuning, distillation, and adaptation can be used to adapt pre-trained models or data. Utilizing pre-training techniques is particularly beneficial when a large amount of pre-training data or resources are available.
V Application Perspectives of Graph Transformers
Graph transformers are finding applications in various domains that involve interconnected data. This section delves into their applications for graph-related tasks, categorized by the level of analysis: node-level, edge-level and graph-level. Beyond these core tasks, graph transformers are also making strides in applications that handle text, images, and videos, where data can be effectively represented as graphs for analysis.
V-A Node-level Tasks
Node-level tasks involve the acquisition of node representations or the prediction of node attributes using the graph structure and node features [153].
V-A1 Protein Structure Prediction
In the field of bioinformatics, graph transformers have demonstrated substantial potential in Protein Structure Prediction (PSP) [154]. Gu et al. [155] introduced HEAL, which employs hierarchical graph transformers on super-nodes that imitate functional motifs to interact with nodes in the protein graph, effectively capturing structural semantics. Pepe et al. [156] used Geometric Algebra (GA) modelling to introduce a new metric based on the relative orientations of amino acid residues, serving as an additional input feature to a graph transformer, assisting in the prediction of the 3D coordinates of a protein. Chen et al. [157] proposed gated-graph transformers integrating node and edge gates within a graph transformer framework to regulate information flow during graph message passing, proving beneficial in predicting the quality of 3D protein complex structures. Despite the encouraging outcomes, various challenges persist, such as the complexity of protein structures, scarcity of high-quality training data, and substantial computational resource [158]. Further investigation is necessary to address these challenges and enhance the precision of these models.
V-A2 Entity Resolution
Entity Resolution (ER) is a crucial task in data management that aims to identify and link disparate representations of real-world entities from diverse sources [159]. Recent research has highlighted the efficacy of graph transformers in ER. For example, Yao et al. [160] proposed Hierarchical Graph Attention Networks (HierGAT) integrating the self-attention mechanism and graph attention network mechanism to capture and leverage the relationships between different ER decisions, leading to substantial enhancements over conventional approaches. Ying et al. [161] extended the standard transformer architecture and introduced several straightforward yet powerful structural encoding techniques to enhance the modeling of graph-structured data. Despite facing challenges related to data complexity and structural information encoding, these techniques have exhibited promising outcomes in terms of enhanced performance, scalability, and accuracy. Furthermore, Dou et al.[162] proposed the Hybrid Matching Knowledge for Entity Matching (GTA) method improves the transformer for representing relational data by integrating additional hybrid matching knowledge acquired through graph contrastive learning on a specially designed hybrid matching graph. This approach has also demonstrated promising results by effectively boosting the transformer for representing relational data and surpassing existing entity matching frameworks.
V-A3 Anomaly Detection
Graph transformers are valuable tools for anomaly detection, especially in dynamic graphs and time series data [65, 163]. They tackle key challenges like encoding information for unattributed nodes and extracting discriminative knowledge from spatial-temporal dynamic graphs. Liu et al [164] proposed TADDY, a transformer-based Anomaly Detection framework, enhancing node encoding to represent each node’s structural and temporal roles in evolving graph streams. Similarly, Xu et al. [165] proposed the Anomaly Transformer which uses an Anomaly-Attention mechanism to measure association discrepancy and employs a minimax strategy to enhance normal-abnormal differentiation. Chen et al. [166] proposed the GTA framework for multivariate time series anomaly detection incorporates graph structure learning, graph convolution, and temporal dependency modeling with a transformer-based architecture. Tuli et al. [167] developed TranAD, a deep transformer network for anomaly detection in multivariate time series data, showing efficient anomaly detection and diagnosis in modern industrial applications. Despite their effectiveness, further research is needed to enhance their performance and applicability across different domains.
V-B Edge-level Tasks
Edge-level tasks aim to learn edge representations or predict edge attributes based on graph structure and node features [168, 169].
V-B1 Drug-Drug Interaction Prediction
Graph transformers have been increasingly employed in the prediction of Drug-Drug Interactions (DDIs) owing to their capability to adeptly model the intricate relationships between drugs and targets [170]. Wang et al. [109] proposed a method which uses a line graph with drug-protein pairs as vertices and a graph transformer network (DTI-GTN) for the purpose of forecasting drug-target interactions. Djeddi et al. [171] proposed a novel approach named DTIOG for the prediction of DTIs, leveraging a Knowledge Graph Embedding (KGE) strategy and integrating contextual information derived from protein sequences. Despite the encouraging outcomes, these approaches encounter challenges such as overlooking certain facets of the intermolecular information and identifying potential interactions for newly discovered drugs [172]. Nevertheless, findings from multiple studies indicate that graph transformers can proficiently anticipate DDIs and surpass the performance of existing algorithms [173].
V-B2 Knowledge Graph Completion
In the domain of Knowledge Graph (KG) completion, the utilization of graph transformers has been extensively investigated [174]. Chen et al. [175] proposed a novel inductive KG representation model, known as iHT, for KG completion through large-scale pre-training. This model comprises an entity encoder and a neighbor-aware relational scoring function, both parameterized by transformers. The application of this approach has led to remarkable advancements in performance, with a relative enhancement of more than 25% in mean reciprocal rank compared to previous models. Liu et al. [176] introduced a generative transformer with knowledge-guided decoding for academic KG completion, which incorporates pertinent knowledge from the training corpus to provide guidance. Chen et al. [177] developed a hybrid transformer with multi-level fusion to tackle challenges in multimodal KG completion tasks. This model integrates visual and textual representations through coarse-grained prefix-guided interaction and fine-grained correlation-aware fusion modules.
V-B3 Recommender Systems
Graph transformers have been effectively utilized in recommender systems by combining generative self-supervised learning with a graph transformer architecture [178]. Xia et al. [179] used the generative self-supervised learning method to extract representations from the data in an unsupervised manner and utilized graph transformer architecture to capture intricate relationships between users and items in the recommendation system. Li et al. [180] introduced a new method for recommender systems that leverages graph transformers (GFormer). Their approach automates the self-supervision augmentation process through a technique called rationale-aware generative self-supervised learning. This technique identifies informative patterns in user-item interactions. The proposed recommender system utilizes a special type of collaborative rationale discovery to selectively augment the self-supervision while preserving the overall relationships between users and items. The rationale-aware self-supervised learning in the graph transformer enables graph collaborative filtering. While challenges remain in areas like graph construction, network design, model optimization, computation efficiency, and handling diverse user behaviors, experiments show that the approach consistently outperforms baseline models on various datasets.
V-C Graph-level Tasks
Graph-level tasks aim to learn graph representations or predict graph attributes based on graph structure and node features.
V-C1 Molecular Property Prediction
Graph transformers are powerful tools for molecular property prediction, utilizing the graph structure of molecules to capture essential structural and semantic information [181]. Chen et al. [182] proposed Algebraic Graph-Assisted Bidirectional Transformer (AGBT) framework which integrates complementary 3D molecular information into graph invariants, rectifying the oversight of three-dimensional stereochemical information in certain machine learning models. Li et al [144] utilized Knowledge-Guided Pre-training of Graph Transformer (KPGT) in a self-supervised learning framework which emphasizes the importance of chemical bonds and models the structural information of molecular graphs. Buterez et al. [183] proposed transfer learning with graph transformer to enhance molecular property prediction on sparse and costly high-fidelity data.
V-C2 Graph Clustering
Graph transformers have been increasingly utilized in the field of Graph Clustering, offering innovative methodologies and overcoming significant challenges [184]. Yun et al. [185] proposed a graph transformer network to generate new graph structures for identifying useful connections between unconnected nodes on the original graph while learning effective node representation on the new graphs in an end-to-end fashion. Gao et al. [186] proposed a patch graph transformer (PatchGT) that segments a graph into patches based on spectral clustering without any trainable parameters and allows the model to first use GNN layers to learn patch-level representations and then use transformer to obtain graph-level representations. These methodologies have addressed issues such as the limitations of the local attention mechanism and difficulties in learning high-level information, leading to enhanced graph representation, improved model performance, and effective node representation.
V-C3 Graph Synthesis
Graph transformers have been applied in the field of graph synthesis to improve graph data mining and representation learning. Existing graph transformers with Positional Encodings have limitations in node classification tasks on complex graphs, as they do not fully capture the local node properties [16]. To address this, Ma et al. [187] introduced the Adaptive Graph Transformer (AGT). This model tackles the challenge of extracting structural patterns from graphs in a way that is both effective and efficient. AGT achieves this by learning from two different graph perspectives: centrality and subgraph views. This approach has been shown to achieve state-of-the-art performance on real-world web graphs and synthetic graphs characterized by heterophily and noise. Jiang et al. [188] proposed an Anchor Graph Transformer (AGFormer) that leverages an anchor graph model to perform more efficient and robust node-to-node message passing for overcoming the computational cost and sensitivity to graph noises of regular graph transformers. Zhu et al. [122] developed a Hierarchical Scalable Graph Transformer (HSGT) which scales the transformer architecture to node representation learning tasks on large-scale graphs by utilizing graph hierarchies and sampling-based training methods.
V-D Other Application Scenarios
Graph transformers have a wide range of applications beyond graph-structured data. They can also be utilized in scenarios involving text, images, or videos. [189].
V-D1 Text Summarization
Text summarization is a crucial aspect of NLP which has been significantly advanced with the introduction of Graph transformers [190]. These models utilize extractive, abstractive, and hybrid methods for summarization. Extractive summarization selects and extracts key sentences or phrases from the original text to create a summary [191]. In contrast, abstractive summarization interprets the core concepts in the text and produces a concise summary. Hybrid summarization combines the advantages of both approaches [192]. Despite the progress, challenges remain in text comprehension, main idea identification, and coherent summary generation. Nevertheless, the application of graph transformers in text summarization has demonstrated promising outcomes in terms of summary quality and efficiency [193].
V-D2 Image Captioning
Graph transformers have emerged as a potent tool within the domain of image captioning, offering a structured representation of images and efficiently processing them to produce descriptive captions [194]. Techniques such as Transforming Scene Graphs (TSG) leverage multi-head attention to architect graph neural networks for embedding scene graphs, which encapsulate a myriad of specific knowledge to facilitate the generation of words across various parts of speech [195]. Despite encountering challenges, such as training complexity, absence of contextual information, and lack of fine-grained details in the extracted features, graph transformers have exhibited promising outcomes. They have enhanced the quality of generated sentences and attained state-of-the-art performance in image captioning endeavors [196].
V-D3 Image Generation
Graph transformers have been effectively utilized in image generation, as demonstrated by various research studies [197]. Sortino et al. [33] proposed a transformer-based method conditioned by scene graphs for image generation, employing a decoder to sequentially compose images. Zhang et al. [198] proposed StyleSwin which uses transformers in constructing a generative adversarial network for high-resolution image creation. Despite challenges like redundant interactions and the requirement for intricate architectures, these studies have exhibited promising outcomes in terms of image quality and variety [199, 200].
V-D4 Video Generation
Graph transformers have been extensively applied in the field of Video Generation. Xiao et al. [201] proposed a Video Graph Transformer (VGT) model, which utilizes a dynamic graph transformer module to encode videos, capturing visual objects, their relationships, and dynamics. It incorporates disentangled video and text transformers for comparing relevance between the video and text. Wu et al. [202] proposed The Object-Centric Video Transformer (OCVT) which adopts an object-centric strategy to break down scenes into tokens suitable for a generative video transformer and understanding the intricate spatiotemporal dynamics of multiple interacting objects within a scene. Yan et al. [203] developed VideoGPT, which learns downsampled discrete latent representations of a raw video through 3D convolutions and axial self-attention. A GPT-like architecture is then used to model the discrete latent in a spatiotemporal manner using position encodings. Tulyakov et al. [204] proposed the MoCoGAN model which creates a video by mapping a series of random vectors to a sequence of video frames. Despite the challenges in capturing complex spatio-temporal dynamics in videos, these methodologies have exhibited promising outcomes across various facets of video generation, ranging from question answering to video summarization and beyond.
VI Open Issues and Future Directions
Despite their immense potential for learning from graph-structured data, graph transformers still face open issues and challenges that require further exploration. Here we highlight some of these open challenges.
VI-A Scalability and Efficiency
The scalability and efficiency of graph transformers pose considerable challenges due to their substantial memory and computational requirements especially when employing global attention mechanisms to deal with large-scale graphs [205, 114]. These challenges are further amplified in deep architectures which are susceptible to overfitting and over-smoothing. To address these issues, several potential strategies can be proposed:
  1. 1.
    Developing efficient attention mechanisms, such as linear, sparse and low-rank attention, to reduce the complexity and memory usage of graph transformers.
  2. 2.
    Applying graph sparsification or coarsening techniques to decrease the size and density of graphs while maintaining their crucial structural and semantic information.
  3. 3.
    Using graph partitioning or sampling methods to divide large graphs into smaller subgraphs or batches for parallel or sequential processing.
  4. 4.
    Exploring graph distillation or compression methods to create compact and effective graph transformer models, which are suitable for deployment on resource-limited devices.
  5. 5.
    Investigating regularization or normalization techniques, such as dropout, graph diffusion, convolution, and graph spectral normalization, to prevent overfitting and over-smoothing in graph transformer models.
VI-B Generalization and Robustness
Graph transformers often face challenges when it comes to generalizing to graphs that they have not encountered before or that fall outside of their usual distribution. This is especially true for graphs that have different sizes, structures, features, and domains. Additionally, graph transformers can be vulnerable to adversarial attacks and noisy inputs, which can result in a decline in performance and the production of misleading results. In order to improve the generalization and robustness of graph transformers, the following strategies could be taken into account:
  1. 1.
    Developing adaptive and flexible attention mechanisms, such as dynamic attention, span-adaptive attention, and multi-head attention with different scales, to accommodate varying graphs and tasks.
  2. 2.
    Applying domain adaptation or transfer learning techniques to facilitate learning from multiple source domains and transfer the knowledge from source domains to target domains.
  3. 3.
    Exploring meta-learning or few-shot learning techniques to enable learning from limited data and rapid adaptation to new tasks.
  4. 4.
    Designing robust and secure attention mechanisms, such as adversarial attention regularization, attention masking, and attention perturbation, to resist adversarial attacks and noisy inputs.
  5. 5.
    Evaluating the uncertainty and reliability of graph transformer models using probabilistic or Bayesian methods, such as variational inference, Monte Carlo dropout, and deep ensembles.
VI-C Interpretability and Explainability
Graph transformers commonly regarded as black box models present significant challenges in terms of interpretability and explainability. The lack of sufficient justification and evidence for their decisions can undermine their credibility and transparency. To address this issue, several approaches can be considered:
  1. 1.
    Developing transparent and interpretable attention mechanisms, such as attention visualization, attention attribution, and attention pruning, to highlight the importance and relevance of different nodes and edges in graphs.
  2. 2.
    Applying explainable artificial intelligence (XAI) techniques, such as saliency maps, influence functions, and counterfactual explanations, to analyze and understand the behaviour and logic of graph transformer models.
  3. 3.
    Exploring natural language generation techniques, such as template-based generation, neural text generation, and question-answering generation, to produce natural language explanations for outputs or actions of graph transformer models.
  4. 4.
    Investigating human-in-the-loop methods, such as active learning, interactive learning, and user studies, to incorporate human guidance in the learning or evaluation process of graph transformer models [31].
VI-D Learning on Dynamic Graphs
Graphs, which are frequently characterized by their dynamic and intricate nature, possess the ability to transform over time as a result of the addition or removal of nodes and edges, as well as the modification of node and edge attributes [206]. Moreover, these graphs may exhibit diverse types and modalities of nodes and edges. In order to empower graph transformers to effectively manage such dynamic graphs, it is advisable to explore the following strategies:
  1. 1.
    Developing temporal and causal attention mechanisms, such as recurrent, temporal, and causal attention, to capture the temporal and causal evolution of graphs.
  2. 2.
    Applying continual learning techniques on dynamic graphs to void forgetting previous knowledge and retraining.
  3. 3.
    Exploring multimodal attention mechanisms, such as image-text, audio-visual, and heterogeneous attention, to integrate multimodal nodes and edges.
  4. 4.
    Leveraging multi-level and multi-layer attention mechanisms, such as node-edge, graph-graph, and hypergraph attention, to aggregate information from different levels and layers of graphs.
VI-E Data Quality and Diversity
Graph transformers require a significant quantity of diverse and high-quality data to achieve effective learning. Nevertheless, in the real world, data is frequently limited, noisy, incomplete, imbalanced, and biased. This has a detrimental effect on the performance and fairness of graph transformer models. To mitigate these challenges related to data quality and diversity, several potential strategies can be considered:
  1. 1.
    Developing noise-tolerant attention mechanisms, such as denoising attention, error correction attention, and self-supervised attention, to filter out data noise and errors.
  2. 2.
    Applying data augmentation or generation techniques, such as graph augmentation [207], graph completion, and graph generation, to enhance the quantity and variety of data.
  3. 3.
    Exploring data imbalance and bias mitigation techniques, such as data re-sampling, data re-weighting, and data debiasing, to improve data fairness and equity.
  4. 4.
    Evaluating data quality and diversity using data quality assessment or data diversity analysis methods, such as data quality metrics, data diversity measures, and data quality visualization.
VII Conclusion
Graph transformers are a novel and powerful class of neural network models, which can effectively encode and process graph-structured data. This survey provides a comprehensive overview of graph transformers in terms of design perspectives, taxonomy, applications, and open issues. We first discuss how graph transformers incorporate graph inductive bias, including node positional encodings, edge structural encodings, message-passing bias and attention bias, to encode the structural information of graphs. Then, we introduce the design of graph attention mechanisms, including global and local attention mechanisms. Afterwards, a taxonomy of graph transformers is presented. This survey also includes a design guide for effective graph transformers, including the best practices and recommendations for selecting appropriate components and hyperparameters. Moreover, the application scenarios of graph transformers are reviewed based on various graph-related tasks (e.g., node-level, edge-level, and graph-level tasks), as well as tasks in other domains. Lastly, current challenges and future directions of graph transformers are identified. This survey aims to serve as a valuable reference for researchers and practitioners interested in graph transformers and their applications.
Acknowledgment
The authors would like to thank Estrid He (RMIT University) for her helpful comments.

### v2

@@
 Finally, we identify remaining challenges in the field, such as scalability and efficiency, generalization and robustness, interpretability and explainability, dynamic and complex graphs, as well as data quality and diversity, charting future directions for graph transformer research.
-Index Terms:
-Graph transformer, attention, graph neural network, representation learning, graph learning, network embedding
+publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
 I Introduction
-Graphs, as data structures with high expressiveness, are widely used to present complex data in diverse domains, such as social media, knowledge graphs, biology, chemistry, and transportation networks [1].
+Graphs, as data structures with high expressiveness, are widely used to present complex data in various domains, such as social networks, knowledge graphs, biology, chemistry, and transportation networks [1].
 They capture both structural and semantic information from data, facilitating various tasks, such as recommendation [2], question answering [3], anomaly detection [4], sentiment analysis [5], text generation [6], and information retrieval [7].
 To effectively deal with graph-structured data, researchers have developed various graph learning models, such as graph neural networks (GNNs), learning meaningful representations of nodes, edges and graphs [8].
-Particularly, GNNs following the message-passing framework iteratively aggregate neighboring information and update node representations, leading to impressive performance on various graph-based tasks [9].
+In particular, GNNs following the message-passing framework iteratively aggregate neighboring information and update node representations, leading to impressive performance on various graph-based tasks [9].
 Applications ranging from information extraction to recommender systems have benefited from GNN modelling of knowledge graphs [10].
-More recently, the graph transformer, as a newly arisen and potent graph learning method, has attracted great attention in both academic and industrial communities [11, 12].
+More recently, the graph transformer, as a new and powerful graph learning method, has attracted great attention in both the academic and industrial communities [11, 12].
 Graph transformer research is inspired by the success of transformers in natural language processing (NLP) [13] and computer vision (CV) [14], coupled with the demonstrated value of GNNs.
+While GNNs rely on local message passing, which limits their ability to capture long-range dependencies and global structure, graph transformers overcome this limitation by using self-attention mechanisms that model interactions across the entire graph.
 Graph transformers incorporate graph inductive bias (e.g., prior knowledge or assumptions about graph properties) to effectively process graph data [15].
-Furthermore, they can adapt to dynamic and heterogeneous graphs, leveraging both node and edge features and attributes.
-[16].
+Furthermore, they can adapt to dynamic and heterogeneous graphs, leveraging both node and edge features and attributes [16].
+However, this expressiveness comes at the cost of higher computational complexity and challenges in scaling to large graphs.
+Thus, a key trade-off in using graph transformers lies between modeling power and computational efficiency.
 Various adaptations and expansions of graph transformers have shown their superiority in tackling diverse challenges of graph learning, such as large-scale graph processing [17].
@@
 Lastly, Min et al.
-[19] concentrated on the architectural design aspect of graph transformers, offering a systematic evaluation of different components on different graph benchmarks, but they did not include significant applications of graph transformers or discuss open issues in this field.
+[19] concentrated on the architectural design aspects of graph transformers, offering a systematic evaluation of different components on different graph benchmarks, but did not include significant applications of graph transformers or discuss open issues in this field.
 To fill these gaps, this survey aims to present a comprehensive and systematic review of recent advancements and challenges in graph transformer research from both design and application perspectives.
@@
     We present a novel taxonomy of graph transformers based on their depth, scalability, and pre-training strategy.
-We also provide a guide to choosing effective graph transformer architectures for different tasks and scenarios.
+We also provide a guide for choosing effective graph transformer architectures for different tasks and scenarios.
   3. 3.
@@
 An overview of this paper is depicted in Figure 1.
-The subsequent survey is structured as follows: Section II introduces notations and preliminaries pertaining to graphs and transformers.
+The subsequent survey is structured as follows: Section II introduces notation and preliminaries pertaining to graphs and transformers.
 Section III delves into the design perspectives of graph transformers that encompass graph inductive bias and graph attention mechanisms.
 Section IV presents a taxonomy of graph transformers categorizing them based on their depth, scalability and pre-training strategy.
-Additionally, a guide is provided for selecting appropriate graph transformer models for diverse tasks and domains.
-Section V explores the application perspectives of graph transformers on various node-level, edge-level, and graph-level tasks, along with other application scenarios.
-Section VI identifies open issues and future directions for research on graph transformers.
-Lastly, Section VII concludes the paper and highlights its main contributions.
+In addition, a guide is provided to select appropriate graph transformer models for various tasks and domains.
+Section V explores the application perspectives of graph transformers in various node-level, edge-level, and graph-level tasks, along with other application scenarios.
+Section VI provides a guide for designing effective graph transformer architectures for different tasks and scenarios.
+Section VII identifies open issues and future directions for research on graph transformers.
+Lastly, Section VIII concludes the paper and highlights its main contributions.
 Figure 1: Organization of this paper
 II Notations and Preliminaries
-In this section, we present fundamental notations and concepts utilized throughout this survey paper.
-Additionally, we provide a concise summary of current methods for graph learning and self-attention mechanisms which serve as the basis for graph transformers.
-Table I includes mathematical notations used in this paper.
+In this section, we present the fundamental notations and concepts used throughout the survey.
+Additionally, we summarize current methods for graph learning and self-attention mechanisms, which form the foundation for graph transformers.
+The notations listed in Table I provide the mathematical language used in subsequent sections, including definitions of graphs, node and edge features, attention parameters, and positional encodings.
+These elements are essential to describe core mechanisms of graph neural networks (Section III), attention computation in graph transformers (Section IV), and task-specific architectures (Section V).
 TABLE I: Notations in this paper
@@
 $G=(V,E)$ A graph with node set $V$ and edge set $E$
-$N$ Number of nodes
+NN Number of nodes
 $M$ Number of edges
@@
 $f$ A message function for the node and edge states
-[…]
+$\textbf{Q},\textbf{K},\textbf{V}$ Query, key and value matrices for self-attention
 $d_{k}$ Dimension of query and key matrices
@@
 II-A Graphs and Graph Neural Networks
-A graph is a data structure consisting of a set of nodes (or vertices) $V$ and a set of edges (or links) $E$ that connect pairs of nodes.
-Formally, a graph can be defined as $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is node set with $N$ nodes and $E=\{e_{1},e_{2},\dots,e_{M}\}$ is edge set with $M$ edges.
-Edge $e_{k}=(v_{i},v_{j})$ indicates the connection between node $v_{i}$ and node $v_{j}$, where $i,j\in\{1,2,\dots,N\}$ and $k\in\{1,2,\dots,M\}$.
-A graph can be represented by an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}$ indicates the presence or absence of an edge between node $v_{i}$ and node $v_{j}$.
-Alternatively, a graph can be represented by the edge list $E\in\mathbb{R}^{M\times 2}$, where each row of $E$ contains the indices of two nodes connected by an edge.
-A graph can also have node features and edge features that describe the attributes or properties of nodes and edges, respectively.
-The features of the nodes can be represented by a feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{n}}$, where $d_{n}$ is the dimension of the node features.
-The edge features can be represented by a feature tensor $\textbf{F}\in\mathbb{R}^{M\times d_{e}}$, where $d_{e}$ is the dimension of the edge features [20].
-Graph learning refers to the task of acquiring low-dimensional vector representations, also known as embeddings for nodes, edges, or the entire graph.
-These embeddings are designed to capture both structural and semantic information of the graph.
-GNNs are a type of neural network model that excels at learning from graph-structured data.
-They achieve this by propagating information along edges and aggregating information from neighboring nodes [21].
-GNNs can be categorized into two main groups: spectral methods and spatial methods.
-Spectral methods are based on graph signal processing and graph Fourier transform, implementing convolution operations on graphs in the spectral domain [22].
-The Fourier graph transform is defined as $\mathbf{\hat{X}}=\textbf{U}^{T}\textbf{X}\textbf{U}$, where $\mathbf{\hat{X}}$ is the spectral representation of node feature matrix X and U is the eigenvector matrix of the normalized graph Laplacian matrix $\textbf{L}=\textbf{I}_{N}-\textbf{D}^{-1/2}\textbf{A}\textbf{D}^{-1/2}$, where $\textbf{I}_{N}$ is the identity matrix and D is diagonal degree matrix with $D_{ii}=\sum_{j=1}^{N}A_{ij}$.
-Spectral methods can capture global information about the graph, but they suffer from high computational complexity, poor scalability, and the lack of generalization to unseen graphs [23].
+A graph is a data structure consisting of a set of nodes (or vertices) $V$ and a set of edges (or links) $E$ connecting pairs of nodes.
+Formally, a graph is defined as $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of nodes with $N$ nodes and $E=\{e_{1},e_{2},\dots,e_{M}\}$ is the set of edges with $M$ edges.
+The edge $e_{k}=(v_{i},v_{j})$ connects node $v_{i}$ and node $v_{j}$, with $i,j\in\{1,2,\dots,N\}$ and $k\in\{1,2,\dots,M\}$.
+A graph can be represented by an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}$ indicates the presence or absence of an edge between nodes $v_{i}$ and $v_{j}$.
+Alternatively, it can be represented by the edge list $E\in\mathbb{R}^{M\times 2}$, where each row of $E$ contains the indices of two connected nodes.
+Graphs can also include node features and edge features, represented by a node feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{n}}$ (where $d_{n}$ is the dimension of node features) and an edge feature tensor $\textbf{F}\in\mathbb{R}^{M\times d_{e}}$ (where $d_{e}$ is the dimension of edge features) [20].
+Graph learning involves deriving low-dimensional vector representations, or embeddings, for nodes, edges, or entire graphs to capture structural and semantic information.
+For instance, embeddings represent users in social networks by encoding interactions; capture atom types and chemical bonds in molecular graphs for drug discovery; and characterize intersections and roads in navigation systems for efficient route planning.
+Graph neural networks (GNNs) effectively learn from graph-structured data by propagating and aggregating information among neighboring nodes [21].
+GNNs are primarily categorized into spectral and spatial methods.
+Spectral methods utilize graph signal processing to perform convolution operations in the spectral domain using the Fourier graph transform defined as $\mathbf{\hat{X}}=\textbf{U}^{T}\textbf{X}\textbf{U}$, where $\mathbf{U}$ contains eigenvectors of the normalized graph Laplacian $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ [22].
+While spectral methods capture global graph properties, they exhibit high computational complexity and limited scalability [23].
 Spatial methods are based on message-passing and neighborhood aggregation, implementing convolution operations on graphs in the spatial domain [8]. The message-passing framework is defined as:
-$\textbf{h}_{v}^{(l+1)}=\phi\left(\textbf{h}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}f(\textbf{h}_{u}^{(l)},\textbf{h}_{v}^{(l)},\textbf{e}_{uv})\right)\end{aligned},$ (1)
+$\begin{aligned} \textbf{h}_{v}^{(l+1)}=\phi\left(\textbf{h}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}f(\textbf{h}_{u}^{(l)},\textbf{h}_{v}^{(l)},\textbf{e}_{uv})\right)\end{aligned},$ (1)
 where $\textbf{h}_{v}^{(l)}$ is the hidden state of node $v$ at layer $l$, $\phi$ is an update function and $\oplus$ is an aggregation function.
-$\mathcal{N}(v)$ is the set of neighbors of node $v$ and $f$ is a message function that depends on node states and edge features.
-$\textbf{e}_{uv}$ is the feature vector of the edge between nodes $u$ and $v$.
-Spatial methods can capture local information of the graph, but they have limitations in modeling long-range dependencies, complex interactions and heterogeneous structures [25].
+$\mathcal{N}(v)$ is the set of neighbors of node $v$ and $f$ is a message function that depends on the states of the nodes and the edge features.
+$\textbf{e}_{uv}$ is the edge feature vector between nodes $u$ and $v$.
+Spatial methods can capture local information on the graph, but have limitations in modeling long-range dependencies, complex interactions, and heterogeneous structures [24].
 II-B Self-attention and Transformers
-Self-attention is a mechanism that enables a model to learn to focus on pertinent sections of input or output sequences [26].
-It calculates a weighted sum of all elements in a sequence with weights determined by the similarity between each element and a query vector.
-Formally, self-attention is defined as:
-$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}\right)\textbf{V}\end{aligned},$ (2)
+Self-attention is a core mechanism that enables models to relate different parts of the input sequence to each other [25].
+Intuitively, it helps the model decide how much attention to pay to other elements when processing a particular item.
+For example, in a sentence, understanding the meaning of a word often depends on its surrounding words; self-attention allows the model to capture such contextual relationships.
+Formally, self-attention computes a weighted sum of all elements in a sequence, where the weights reflect the similarity between each element and a query vector.
+It is defined as:
+$\begin{aligned} \text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}\right)\textbf{V}\end{aligned},$ (2)
 where Q, K and V are query, key and value matrices, respectively.
 $d_{k}$ is the dimension of the query and key matrices.
-Self-attention can capture long-range dependencies, global context and variable-length sequences without using recurrence or convolution.
-Figure 2: An illustration of transformer architecture [13].
+Self-attention can capture long-range dependencies, global context, and variable-length sequences without using recurrence or convolution.
 Transformers are neural network models that use self-attention as the main building block [13].
@@
 The decoder takes an output sequence $\textbf{Y}=\{\textbf{y}_{1},\textbf{y}_{2},\dots,\textbf{y}_{N}\}$ and generates a sequence of hidden states $\textbf{S}=\{\textbf{s}_{1},\textbf{s}_{2},\dots,\textbf{s}_{N}\}$.
-The decoder also uses an attention mechanism to attend to the encoder's hidden states.
-Formally, the encoder and decoder are defined as:
-$\textbf{z}_{i}$ $=\text{EncoderLayer}(\textbf{x}_{i},\textbf{Z}_{<i}),$ (3)
-$\textbf{s}_{j}$ $=\text{DecoderLayer}(\textbf{y}_{j},\textbf{S}_{<j},\textbf{Z}).$
-Here, EncoderLayer and DecoderLayer are composed of multiple self-attention and feed-forward sublayers.
-Transformers can achieve state-of-the-art results on various tasks, such as machine translation [27], text mining [28, 29], document comprehending [30, 31], image retrieval [32], visual question answering [3], image generation [33], and recommendation systems [34, 35].
-An overview of vanilla transformer is shown in Figure 2.
-Graph transformers integrate graph inductive bias into transformer to acquire knowledge from graph-structured data [36].
+In the final step of the decoder, these hidden states are passed through a softmax layer to compute probabilities over possible output tokens, which can then be used for downstream tasks such as sequence generation or classification.
+The decoder also uses an attention mechanism to attend to the encoder’s hidden states.
+Formally, the encoder and decoder are defined as: $\textbf{z}_{i}=\text{EncoderLayer}(\textbf{x}_{i},\textbf{Z}_{<i})$, $\textbf{s}_{j}=\text{DecoderLayer}(\textbf{y}_{j},\textbf{S}_{<j},\textbf{Z}).$ Here, EncoderLayer and DecoderLayer are composed of multiple self-attention and feed-forward sublayers.
+Transformers can achieve state-of-the-art results in various tasks, such as machine translation [26], text mining [27], document comprehension [28], image retrieval [29], visual question answering [3], and image generation [30].
+An overview of the vanilla transformer is shown in Figure 2.
+Graph transformers integrate graph inductive bias into transformers to acquire knowledge from graph-structured data [31].
 By employing self-attention mechanisms on nodes and edges, graph transformers can effectively capture both local and global information of the graph.
-In particular, graph transformers exhibit the ability to handle heterogeneous graphs containing diverse types of nodes and edges, as well as complex graphs featuring higher-order structures [12, 37].
+However, traditional transformer models, which are designed for sequence data, cannot be directly applied to graph data because graph-structured data does not have a natural ordering, as is the case with sequence data.
+In graph data, nodes and edges are interconnected in a non-sequential manner, and this necessitates the modification of the standard self-attention mechanism.
+To address this challenge, graph transformers modify the attention mechanism by incorporating graph-specific inductive biases such as positional encoding based on graph structure and attention over both node and edge features.
+These adaptations enable graph transformers to learn dependencies not only along the sequential dimensions but also in the graph topology, capturing relationships between nodes even in non-sequential configurations.
+[12, 32].
+Figure 2: An illustration of transformer architecture [13].
 III Design Perspectives of Graph Transformers
-In this section, we discuss the primary architectures of graph transformers, aiming to explore their design perspectives in depth.
-Particularly, we will focus on two key components: graph inductive biases and graph attention mechanisms, to understand how these elements shape graph transformer models’ capabilities.
+In this section, we introduce the concept of “design perspectives” in Graph Transformers, referring to the core principles and strategies that guide how graph-specific inductive biases, attention mechanisms, and structural properties are integrated into the Transformer architecture.
+By examining these underlying design considerations, we aim to illustrate how Graph Transformers can effectively capture both local and global relationships in graph-structured data.
+While this section focuses on fundamental design concepts, Section IV provides a detailed taxonomy of Graph Transformer architectures, covering the diverse ways in which these design principles are operationalized to address various tasks and trade-offs in graph learning.
+This section examines how graph inductive biases and graph attention mechanisms are incorporated into graph transformer models to handle structured data effectively.
 III-A Graph Inductive Bias
-Unlike Euclidean data, such as texts and images, graph data is non-Euclidean data, which has intricate structures and lacks a fixed order and dimensionality, posing difficulties in directly applying standard transformers on graph data [16].
-To address this issue, graph transformers incorporate graph inductive bias to encode the structural information of graphs and achieve effective generalization of transformers across new tasks and domains.
-In this section, we explore the design perspectives of graph transformers through the lens of graph inductive bias.
-We classify graph inductive bias into four categories: node positional bias, edge structural bias, message-passing bias, and attention bias (see Figure.
-3).
-Figure 3: Graph Inductive bias: (A) and (B) are local and global node positional biases, respectively.
-(C) and (D) are local and global edge structural biases, respectively.
-(E) and (F) are local and global attention biases, respectively.
-Notably, message-passing bias is a kind of local attention bias.
+Unlike Euclidean data such as text and images, graph data is non-Euclidean with complex structures and no fixed order or dimensionality, making it challenging to apply standard transformers [16].
+Graph transformers address this issue by integrating graph inductive bias, which encodes structural information for better generalization.
+We classify graph inductive bias in graph transformers into four categories: node positional bias, edge structural bias, message-passing bias, and attention bias.
 III-A1 Node Positional Bias
-Node positional bias is a crucial inductive bias for graph transformers because it provides information about the relative or absolute positions of nodes in a graph [38].
-Formally, given a graph $G=(V,E)$ with $N$ nodes and $M$ edges, each node $v_{i}\in V$ has a feature vector $\textbf{x}_{i}\in\mathbb{R}^{d}_{n}$.
-A graph transformer aims to learn a new feature vector $\textbf{h}_{i}\in\mathbb{R}^{d}_{k}$ for each node by applying a series of self-attention layers.
-A self-attention layer can be defined as:
-$\textbf{h}_{i}=\sum_{j=1}^{n}a_{ij}W\textbf{x}_{j}+b,$ (4)
-where $a_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, measureing the relevance or similarity of their features.
-$W$ and $b$ are learnable parameters.
-However this self-attention mechanism does not consider the structural and positional information of nodes, which is crucial for capturing graph semantics and inductive biases [39].
-Node positional encoding is a way to address this challenge by providing additional positional features to nodes, reflecting their relative or absolute positions in the graph.
-Local Node Positional Encodings.
-Building on the success of relative positional encodings in NLP, graph transformers leverage a similar concept for local node positional encodings.
-In NLP, each token receives a feature vector that captures its relative position and relationship to other words in the sequence [40].
-Likewise, graph transformers assign feature vectors to nodes based on their distance and relationships with other nodes in the graph [41].
-This encoding technique aims to preserve the local connectivity and neighborhood information of nodes, which is critical for tasks like node classification, link prediction, and graph generation.
-A proficient approach for integrating local node positional information involves the utilization of one-hot vectors.
-These vectors represent the hop distance between a node and its neighboring nodes [11]:
-$\textbf{p}_{i}=[I(d(i,j)=1),I(d(i,j)=2),...,I(d(i,j)=max)].$ (5)
-In this equation, $d(i,j)$ represents the shortest path distance between node $v_{i}$ and node $v_{j}$ and $I$ is an indicator function that returns 1 if its argument is true and 0 otherwise.
-The maximum hop distance is denoted by $max$.
-This encoding technique was utilized by Velickovic et al.
-[42] to enhance their Graph Attention Networks (GATs) with relative position-aware self-attention.
-Another way to incorporate local node positional encodings in a graph is by using learnable embeddings that capture the relationship between two nodes [43, 44].
-This approach is particularly useful when a node has multiple neighbors with different edge types or labels.
-In this case, the local node positional encoding can be learned based on these edge features:
-$\textbf{p}_{i}=[f(\textbf{e}_{ij_{1}}),f(\textbf{e}_{ij_{2}}),...,f(\textbf{e}_{ij_{l}})],$ (6)
-where $\textbf{e}_{ij}$ is edge feature between node $v_{i}$ and node $v_{j}$, $f$ is a learnable function that maps edge features to embeddings and $l$ is the number of neighbors considered.
-To enhance the implementation of local node positional encodings in a broader context, a viable strategy is leveraging graph kernels or similarity functions to evaluate the structural similarity between two nodes within the graph [45, 37].
-For instance, when a node exhibits three neighbors with unique subgraph patterns or motifs in their vicinity, its local node positional encoding can be computed as a vector of kernel values between the node and its neighboring nodes:
-$\textbf{p}_{i}=[K(G_{i},G_{j_{1}}),K(G_{i},G_{j_{2}}),...,K(G_{i},G_{j_{l}})].$ (7)
-In this equation, $G_{i}$ refers to the subgraph formed by node $v_{i}$ and its neighbors.
-The function $K$ is a graph kernel function that measures the similarity between two subgraphs.
+Node positional bias provides relative or absolute positional information for nodes in a graph [33].
+Given a graph $G=(V,E)$ with $N$ nodes and $M$ edges, each node $v_{i}\in V$ has a feature vector $\textbf{x}_{i}\in\mathbb{R}^{d}_{n}$.
+A graph transformer aims to learn a new embedding vector $\textbf{h}_{i}\in\mathbb{R}^{d}_{k}$ for each node by leveraging self-attention layers during training.
+A self-attention layer is defined as: $\textbf{h}_{i}=\sum_{j=1}^{n}a_{ij}W\textbf{x}_{j}+b,$ where $a_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, measuring the relevance or similarity of their features, and $W$ and $b$ are learnable parameters.
+Because this mechanism initially lacks structural and positional information from the nodes, node positional encoding addresses these gaps by adding positional features that capture graph semantics and inductive biases [34].
+Local Node Positional encoding.
+Building on the success of relative positional encoding in NLP, graph transformers adopt a similar concept for local node positional encoding.
+In NLP, each token receives a feature vector capturing its relative position and relationship to other words [35].
+Similarly, graph transformers assign feature vectors to nodes based on their distance and relationships with other nodes [36], preserving local connectivity and neighborhood information, which benefits tasks such as node classification, link prediction, and graph generation.
+A frequently used approach leverages one-hot vectors representing the hop distance between a node and its neighbors [11]: $\textbf{p}_{i}=[I(d(i,j)=1),I(d(i,j)=2),...,I(d(i,j)=max)],$ where $d(i,j)$ is the shortest path distance between nodes $v_{i}$ and $v_{j}$, and $I$ is an indicator function.
+The maximum hop distance is maxmax.
+This technique was used by Velickovic et al.
+[37] to enhance Graph Attention Networks (GATs) with position-aware self-attention.
+Another method involves learnable embeddings capturing the relationship between nodes [38, 39], useful when a node has multiple neighbors with different edge types or labels: $\textbf{p}_{i}=[f(\textbf{e}_{ij_{1}}),f(\textbf{e}_{ij_{2}}),...,f(\textbf{e}_{ij_{l}})],$ where $\textbf{e}_{ij}$ is the edge feature between nodes $v_{i}$ and $v_{j}$, $f$ is a learnable function mapping edge features to embeddings, and $l$ is the number of neighbors considered.
+Another option is to leverage graph kernels or similarity functions to measure structural similarity between nodes [40, 32]: $\textbf{p}_{i}=[K(G_{i},G_{j_{1}}),K(G_{i},G_{j_{2}}),...,K(G_{i},G_{j_{l}})],$ where $G_{i}$ is the subgraph consisting of node $v_{i}$ and its neighbors, and $K$ is a graph kernel function measuring subgraph similarity.
 Mialon et al.
-[46] utilized this approach in their GraphiT model, which incorporates positive definite kernels on graphs as relative positional encodings for graph transformers.
-Local node positional encodings enhance the self-attention mechanism of graph transformers by integrating structural and topological information of nodes.
-This encoding approach offers the advantage of preserving the sparsity and locality of graph structure, resulting in enhanced efficiency and interpretability.
-However, a limitation of this method is its restricted ability to capture long-range dependencies or global properties of graphs, which are essential for tasks such as graph matching or alignment.
-Global Node Positional Encodings.
-The concept of global node positional encodings is inspired by the use of absolute positional encodings in NLP [47].
-In NLP, as mentioned previously, every token receives a feature vector indicating its position within a sequence.
-Extending this idea to graph transformers, each node can be assigned a feature vector representing its position within the embedding space of the graph [48].
-The objective of this encoding technique is to encapsulate the overall geometry and spectrum of graphs, thereby unveiling its intrinsic properties and characteristics.
-One method for obtaining global node positional encodings is to leverage eigenvectors or eigenvalues of a matrix representation, such as adjacency matrix or Laplacian matrix [39].
-For instance, if a node’s coordinates lie within the first k eigenvectors of graph Laplacian, its global node positional encoding can be represented by the coordinate vector:
-$\textbf{p}_{i}=[u_{i1},u_{i2},...,u_{ik}],$ (8)
-where $u_{ij}$ is $j$-th component of $i$-th eigenvector of the graph Laplacian matrix.
-One alternative approach to incorporating global node positional encodings is by utilizing diffusion or random walk techniques, such as personalized PageRank or heat kernel [49].
-For example, if a node possesses a probability distribution over all other nodes in the graph, following a random walk [50], its global node positional encoding can be represented by this probability vector:
-$\textbf{p}_{i}=[\pi_{i1},\pi_{i2},...,\pi_{iN}],$ (9)
-where $\pi_{ij}$ is the probability of reaching node $v_{j}$ from node $v_{i}$ after performing a random walk on the graph.
-A more prevalent approach to implementing global node positional encodings is to utilize graph embedding or dimensionality reduction techniques that map nodes to a lower-dimensional space while maintaining a sense of similarity or distance [51].
-For instance, if a node possesses coordinates in a space derived from applying multi-dimensional scaling or graph neural networks to the graph, its global node positional encoding can be represented by that coordinate vector:
-$\textbf{p}_{i}=[y_{i1},y_{i2},...,y_{ik}],$ (10)
-where $y_{ij}$ is $j$-th component of $i$-th node embedding in a $k$-dimensional space, which can be obtained by minimizing the objective function that preserves graph structure:
-$\min_{\textbf{Y}}\sum_{i,j=1}^{N}\textbf{w}_{ij}\|\textbf{y}_{i}-\textbf{y}_{j}\|^{2}.$ (11)
-Here, $\textbf{w}_{ij}$ is a weight matrix that reflects the similarity or distance between node $v_{i}$ and node $v_{j}$, $Y$ is the matrix of node embeddings.
-The primary aim of global node positional encodings is to improve the representation of node attributes in graph transformers by incorporating geometric and spectral information from graphs.
-This encoding method offers the advantage of capturing long-range dependencies and overall graph characteristics, benefiting tasks like graph matching and alignment.
-However, a drawback of this encoding approach is that it could undermine the sparsity and locality of graph structures, potentially impacting efficiency and interpretability.
+[41] used this idea in their GraphiT model, applying positive definite kernels as relative positional encoding.
+Although local node positional encoding preserves the sparsity and locality of graph structure and can improve efficiency and interpretability, it has limited ability to capture long-range dependencies or global properties of graphs, which are often required for tasks like graph matching or alignment.
+Global Node Positional encoding.
+Global node positional encoding takes inspiration from absolute positional encoding in NLP [42], assigning to each node a characteristic vector that represents its position within the embedding space of the graph [43].
+One method for obtaining global node positional encoding leverages eigenvectors or eigenvalues of a matrix representation (e.g., adjacency or Laplacian) [34]: $\textbf{p}_{i}=[u_{i1},u_{i2},...,u_{ik}],$ where $u_{ij}$ is the $j$-th component of the $i$-th eigenvector of the graph Laplacian.
+An alternative approach relies on diffusion or random walk techniques, such as personalized PageRank or heat kernel [44].
+In this case, the global positional encoding of a node can be represented by its random walk transition probabilities [45]: $\textbf{p}_{i}=[\pi_{i1},\pi_{i2},...,\pi_{iN}],$ where $\pi_{ij}$ is the probability of reaching node $v_{j}$ from node $v_{i}$.
+Another common method is graph embedding or dimensionality reduction, mapping nodes into a lower-dimensional space while preserving similarity or distance [46].
+In such cases: $\textbf{p}_{i}=[y_{i1},y_{i2},...,y_{ik}],$ where $y_{ij}$ is the $j$-th component of the $i$-th node embedding, obtained by minimizing an objective function that preserves the graph’s structure: $\min_{\textbf{Y}}\sum_{i,j=1}^{N}\textbf{w}_{ij}\|\textbf{y}_{i}-\textbf{y}_{j}\|^{2}.$ Here, $\textbf{w}_{ij}$ is a weight matrix reflecting node similarity or distance.
+The primary aim of global node positional encoding is to enhance node attribute representation by incorporating geometric and spectral graph information.
+This method captures long-range dependencies and overall graph characteristics, aiding tasks like graph matching and alignment.
+However, it can reduce sparsity and locality in graph representations, potentially affecting efficiency and interpretability.
 III-A2 Edge Structural Bias
-In the realm of graph transformers, edge structural bias is crucial for extracting and understanding complex information within graph structure [52].
-Edge structural bias is versatile and can represent various aspects of graph structure, including node distances, edge types, edge directions and local sub-structures.
-Empirical evidence has shown that edge structural encodings can improve the effectiveness of graph transformers [53, 54, 55].
-Local Edge Structural Encodings.
-Local edge structural encodings capture the local structure of a graph by encoding relative position or distance between two nodes [52].
-These encodings borrow ideas from relative positional encodings used in NLP and CV, where they are used for modeling sequential or spatial order of tokens or pixels [56].
-However, in the context of graphs, the concept of relative position or distance between nodes becomes ambiguous due to the presence of multiple connecting paths with varying lengths or weights.
-Consequently, the scientific community has proposed various methods to define and encode this information specifically for graph structures.
-GraphiT [46] introduces local edge structural encodings to graph transformers.
-It promotes the use of positive definite kernels on graphs to measure node similarity considering the shortest path distance between them.
-The kernel function is defined as:
-$k(u,v)=\exp(-\alpha d(u,v)),$ (12)
-where $u$ and $v$ are two nodes in a graph, $d(u,v)$ is their shortest path distance and $\alpha$ is a hyperparameter that controls decay rate.
-The kernel function is then used to modify the self-attention score between two nodes as:
-$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+k(\textbf{Q},\textbf{K}))\textbf{V},$ (13)
-where $k(\textbf{Q},\textbf{K})$ is the matrix of kernel values computed for each pair of nodes.
-EdgeBERT [57] proposes using edge features as additional input tokens for graph transformers.
-These edge features are obtained by applying a learnable function to the source and target node features of each edge.
-The resulting edge features are then concatenated with node features and fed into a standard transformer encoder.
-More recently, the Edge-augmented Graph Transformer (EGT) [58] introduces residual edge channels as a mechanism to directly process and output both structural and node information.
-The residual edge channels are matrices that store edge information for each pair of nodes.
-They are initialized with either an adjacency matrix or the shortest path matrix and updated at each transformer layer by applying residual connections.
-These channels are then used to adjust the self-attention score between two nodes
-$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V},\textbf{R}_{e})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+\textbf{R}_{e})\textbf{V},$ (14)
-where $\textbf{R}_{e}$ is the residual edge channel matrix.
-Although local edge structural encodings can capture detailed structural information, they may have limitations in capturing overall structural information.
-This can lead to increased computational complexity or memory usage as pairwise information needs to be computed and stored.
-Additionally, the effectiveness of these encodings may vary depending on the selection and optimization of encoding or kernel function for different graphs and tasks.
-Global Edge Structural Encodings.
-Global edge structural encodings aim to capture the overall structure of a graph.
-Unlike NLP and CV domains, the exact position of a node in graphs is not well-defined because there is no natural order or coordinate system [59].
-Several approaches have been suggested to tackle this issue.
-GPT-GNN [60] is an early work that utilizes graph pooling and unpooling operations to encode the hierarchical structure of a graph.
-It reduces graph size by grouping similar nodes and then restoring the original size by assigning cluster features to individual nodes.
-This approach results in a multiscale representation of graphs and has demonstrated enhanced performance on diverse tasks.
-Graphormer [11] uses spectral graph theory to encode global structure.
-It uses eigenvectors of normalized Laplacian matrix as global positional encodings for nodes.
-This method can capture global spectral features (e.g., connectivity, centrality and community structure).
+In graph transformers, edge structural bias provides a way to extract and interpret complex graph structure information [47].
+It can capture diverse structural aspects, including node distances, edge types, edge directions, and local sub-structures.
+Studies indicate that edge structural encoding can improve the performance of graph transformers [48, 49, 50].
+Local Edge Structural encoding.
+Local edge structural encoding captures the local structure of a graph by encoding the relative position or distance between nodes [47].
+Drawing on relative positional encoding in NLP and CV [51], these methods must account for the fact that graph distances can be ambiguous due to multiple paths with varying lengths or weights.
+GraphiT [41] introduces local edge structural encoding by using positive definite kernels to measure node similarity based on shortest path distance: $k(u,v)=\exp(-\alpha d(u,v)),$ where $u$ and $v$ are nodes, $d(u,v)$ is the shortest path distance, and $\alpha$ controls the decay rate.
+This kernel modifies the self-attention score: $\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+k(\textbf{Q},\textbf{K})\right)\textbf{V},$ where $k(\textbf{Q},\textbf{K})$ is the matrix of kernel values.
+EdgeBERT [52] uses edge features as additional input tokens, obtained by applying a learnable function to source and target node features of each edge, then concatenating them with node features for transformer processing.
+The Edge-augmented Graph Transformer (EGT) [53] introduces residual edge channels, matrices that store edge information for each node pair.
+Initialized with an adjacency matrix or shortest path matrix, these channels are updated at each transformer layer using residual connections to adjust self-attention scores: $\text{Attention}(\textbf{Q},\textbf{K},\textbf{V},\textbf{R}_{e})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+\textbf{R}_{e}\right)\textbf{V},$ where $\textbf{R}_{e}$ is the residual edge channel matrix.
+Local edge structural encoding can capture fine-grained structural information but may overlook global graph properties, and it may increase computational cost depending on the chosen kernel or encoding strategy.
+Global Edge Structural encoding.
+Global edge structural encoding aims to represent the overall structure of a graph, which is difficult due to the absence of a natural order or coordinate system [54].
+GPT-GNN [55] uses graph pooling and unpooling operations to encode hierarchical structures, reducing the graph size by grouping similar nodes and then restoring the original size by assigning cluster features back to individual nodes.
+Graphormer [11] applies spectral graph theory, using eigenvectors of the normalized Laplacian matrix as global positional encoding to capture global spectral features such as connectivity and community structure.
 Park et al.
-[41] extended Graphormer by using singular value decomposition (SVD) to encode global structure.
-They utilized the left singular matrix of the adjacency matrix as global positional encodings for nodes.
-This approach can handle both symmetric and asymmetric matrices.
-Global edge structural encodings excel at capturing coarse-grained structural information at the graph level, benefiting tasks that require global understanding.
-However, they may struggle with capturing fine-grained node-level information and can lose data or introduce noise during encoding.
-In addition, their effectiveness may depend on the choice of encoding technique and matrix representations.
+[36] extended this by using singular value decomposition (SVD), employing the left singular matrix of the adjacency matrix as global positional encoding, suitable for both symmetric and asymmetric matrices.
+Global edge structural encoding effectively captures coarse-grained structure at the graph level but may lose detail about individual nodes or local patterns, and its success depends on the chosen encoding technique and matrix representations.
 III-A3 Message-passing Bias
-Message-passing bias is a crucial inductive bias for graph transformers to facilitate learning from the local structure of graphs [61].
-This bias enables graph transformers to exchange information between nodes and edges, thereby capturing dependencies and interactions among graph elements.
-Moreover, message-passing bias helps graph transformers in overcoming certain limitations of standard transformer architecture, such as the quadratic complexity of self-attention, the absence of positional information and challenges associated with handling sparse graphs [39].
-Formally, message-passing bias can be expressed as follows:
-$\textbf{h}_{v}^{(t+1)}=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$ (15)
-Here, $\textbf{h}_{v}^{(t)}$ is the representation of node $v$ at layer $t$, $u$ is a neighboring node of node $v$, $\textbf{e}_{uv}$ is the feature of the edge between nodes $v$ and $u$, and $f$ is the function that aggregates information from neighbors and edges.
-Incorporation of message-passing bias in graph transformers can be achieved through three primary approaches: preprocessing, interleaving and post-processing (as illustrated in Figure 4 ).
-These approaches vary in how they combine message-passing operations with self-attention layers within transformer architecture.
+Message-passing bias enables graph transformers to utilize local graph structure by exchanging information between nodes and edges, addressing certain limitations of the standard transformer architecture, such as the quadratic complexity of self-attention and the lack of positional information [56, 34].
+Approaches to incorporate this bias in graph transformers include preprocessing, interleaving, and post-processing, each employing distinct methods of combining message-passing operations with self-attention layers (see Figure 3).
 Figure 3: Incorporation of message-passing bias in transformer model.
 Preprocessing.
-Preprocessing involves applying message-passing operations to node features before feeding them to self-attention layers [36].
-This technique aims to augment node features with local structural information, making them more compatible with global self-attention.
-Preprocessing can maintain the integrity of the original transformer architecture without making any modifications and utilize message-passing modules from existing GNNs.
-It can be mathematically defined as follows:
+Preprocessing applies message-passing operations to node features before self-attention layers [31], thereby augmenting node features with local structural information.
+This approach retains the original transformer design and employs standard GNN-based message-passing modules.
+Mathematically, it is defined as:
 $\textbf{h}_{v}^{(t)}$ $=f(\textbf{h}_{v}^{(t-1)},\{\textbf{h}_{u}^{(t-1)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$ (3)
-$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}).$
-Here, $\textbf{h}_{v}^{(0)}=\textbf{x}_{v}$, $\textbf{x}_{v}$ is input feature of node $v$, SelfAttention is a function that performs self-attention over all nodes.
-The limitation of preprocessing is that it applies message-passing only once before self-attention layers.
-This approach may not fully capture intricate interactions between nodes and edges on various scales.
-Additionally, the preprocessing step may introduce redundancy and inconsistency between the message-passing module and the self-attention layer as they both serve similar functions of aggregating information from neighboring elements.
+$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$
+where $\textbf{h}_{v}^{(0)}=\textbf{x}_{v}$.
+Although this method provides local structural cues, applying message-passing only once before self-attention layers may limit capturing multi-scale interactions.
+It can also introduce redundancy between the message-passing module and the self-attention layer.
 Interleaving.
-Interleaving refers to the technique employed in graph transformer architecture that involves alternating message-passing operations and self-attention layers [62, 16, 63].
-The objective of this technique is to achieve a balance between local and global information processing, thereby enabling multi-hop reasoning over graphs.
-By integrating message-passing modules into core components of graph transformers, interleaving enhances their expressive power and flexibility.
-It can be mathematically defined as follows:
+Interleaving alternates message-passing operations and self-attention layers [57, 16, 58], balancing local and global information processing.
+This arrangement supports multi-hop reasoning and can improve the expressive power of graph transformers.
+Mathematically, it is expressed as:
 $\textbf{h}_{v}^{(t+1)}$ $=\theta+\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$ (4)
-$\theta$ $=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
-One drawback of the interleaving technique is its impact on the complexity and computational requirements of graph transformers.
-This is due to the need for additional parameters and operations compared to pre-processing or post-processing.
-Furthermore, interleaving can potentially lead to interference and conflict between the message-passing module and the self-attention layer as they each update node representations in distinct manners.
-Post-processing.
-Post-processing refers to the technique of applying message-passing operations to node representations obtained from self-attention layers [64, 58].
-The purpose of this approach is to refine and adjust node representations based on underlying graph structure, thereby enhancing their interpretability and robustness.
-By doing so, this method aims to improve the quality and utility of node representations for downstream tasks and applications.
-It can be mathematically defined as follows:
+$\theta$ $=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$
+where added parameters and operations can increase computational cost, and conflicting updates by message-passing and self-attention may require careful tuning.
+Post-processing. Post-processing applies message-passing to the node representations produced by self-attention layers [59, 53], refining them based on the underlying graph structure. Mathematically:
 $\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$ (5)
-$\textbf{h}_{v}^{(T+1)}$ $=f(\textbf{h}_{v}^{(T)},\{\textbf{h}_{u}^{(T)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
-Here, $T$ is the final layer of the graph transformer.
-The drawback of post-processing is its limited application of message-passing after self-attention layers, potentially failing to capture intricate semantics and dynamics of graph data.
-Additionally, post-processing runs the risk of introducing noise and distortion to node representations as it has the potential to overwrite or override information acquired by self-attention layers.
+$\textbf{h}_{v}^{(T+1)}$ $=f(\textbf{h}_{v}^{(T)},\{\textbf{h}_{u}^{(T)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$
+where $T$ denotes the final layer of the graph transformer.
+Since message-passing occurs only after self-attention, it may overlook complex interactions at earlier layers.
+Furthermore, overwriting information from self-attention layers can introduce inconsistencies in the final node representations.
 III-A4 Attention Bias
-Attention bias enables graph transformers to effectively incorporate graph structure information into the attention mechanism without message-passing or positional encodings [39, 65].
-Attention bias modifies attention scores between nodes based on their relative positions or distances in the graph.
-It can be categorized as either local or global depending on whether it focuses on the local neighborhood or global topology of the graph.
-Local Attention Bias.
-Local attention bias limits the attention to a local neighborhood surrounding each node.
-This concept is analogous to the message-passing mechanism observed in GNNs [66].
-It can be mathematically defined as follows:
-[…]
-where $\alpha_{ij}$ is the attention score between node $v_{i}$ and node $v_{j}$, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are their node features, $g$ is a function that computes the similarity between two nodes, such as dot-product and linear transformation.
-$b_{ij}$ is a local attention bias term that modifies attention score based on the distance between node $v_{i}$ and node $v_{j}$.
-The local attention bias term can be either a binary mask that only allows attention within a certain hop distance [67, 42, 36] or a decay function that decreases attention score with increasing distance [68, 64].
+Attention bias allows graph transformers to incorporate graph structure information into the attention mechanism without needing message-passing or positional encoding [34, 60].
+It modifies attention scores between nodes based on their relative positions or distances in the graph and can be categorized as either local or global, depending on whether it focuses on the local neighborhood or global topology of the graph.
+Local Attention Bias. Local attention bias restricts attention to a node’s local neighborhood, similar to the message-passing mechanism in GNNs [61]. It is mathematically defined as:
+$\alpha_{ij}=\frac{\exp(g(\mathbf{x}_{i},\mathbf{x}_{j})\cdot b_{ij})}{\sum_{k\in\mathcal{N}(v_{i})}\exp(g(\mathbf{x}_{i},\mathbf{x}_{k})\cdot b_{ik})},$ (6)
+where $\alpha_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are their node features, and $g$ computes node similarity, such as dot-product or linear transformation.
+The term $b_{ij}$ is a local attention bias that adjusts the attention score based on the distance between nodes $v_{i}$ and $v_{j}$.
+This bias can be a binary mask allowing attention within a certain hop distance [62, 37, 31] or a decay function decreasing attention with increasing distance [63, 59].
 Global Attention Bias.
-Global attention bias integrates global topology information into the attention mechanism independent of message-passing and positional encodings [69].
+Global attention bias integrates global topology information into the attention mechanism independent of message-passing and positional encoding [64].
 It can be mathematically defined as follows:
@@
 III-B Graph Attention Mechanisms
-Graph attention mechanisms play an important role in the construction of graph transformers [71].
-By dynamically assigning varying weights to nodes and edges in the graph, these mechanisms enable transformers to prioritize and emphasize the most relevant and important elements for a given task [72].
-Specifically, a graph attention mechanism is a function that maps each node $v_{i}\in V$ to a vector $\textbf{h}_{i}\in\mathbb{R}^{d_{k}}$:
-$\textbf{h}_{i}=f_{n}(\textbf{x}_{i},\{\textbf{x}_{j}\}_{v_{j}\in\mathcal{N}(v_{i})},\{A_{ij}\}_{v_{j}\in\mathcal{N}(v_{i})}),$ (21)
-where $f_{n}$ is a nonlinear transformation, $\textbf{x}_{i}$ is the input feature vector of node $v_{i}$, $\textbf{x}_{j}$ is the input feature vector of node $v_{j}$.
-The function $f_{n}$ can be decomposed into two parts: an attention function and an aggregation function.
-The attention function computes a scalar weight for each neighbor of node $v_{i}$, denoted by $\alpha_{ij}$, which reflects the importance or relevance of node $v_{j}$ for node $v_{i}$:
-$\alpha_{ij}=\text{softmax}_{i}(\text{LeakyReLU}(\textbf{W}_{a}[\textbf{x}_{i}||\textbf{x}_{j}])),$ (22)
-where $\textbf{W}_{a}\in\mathbb{R}^{1\times 2d_{n}}$ is a learnable weight matrix, ||||| | is the concatenation operation, and $\text{softmax}_{i}$normalizes weights over all neighbors of node $v_{i}$.
-The aggregation function combines the weighted features of the neighbors to obtain the output representation of node $v_{i}$:
+Graph attention mechanisms are widely used in constructing graph transformers [66].
+By dynamically weighting nodes and edges, they enable transformers to focus on the most relevant parts of a graph for a given task [67].
+Formally, a graph attention mechanism maps each node $v_{i}\in V$ to a vector $\textbf{h}_{i}\in\mathbb{R}^{d_{k}}$: $\textbf{h}_{i}=f_{n}(\textbf{x}_{i},\{\textbf{x}_{j}\}_{v_{j}\in\mathcal{N}(v_{i})},\{A_{ij}\}_{v_{j}\in\mathcal{N}(v_{i})}),$ where $f_{n}$ is a nonlinear function, $\textbf{x}_{i}$ is the feature of node $v_{i}$, and $\textbf{x}_{j}$ is the feature of neighbor $v_{j}$.
+The mechanism typically comprises an attention function and an aggregation function.
+The attention function computes a scalar weight $\alpha_{ij}$ for each neighbor: $\alpha_{ij}=\text{softmax}_{i}(\text{LeakyReLU}(\textbf{W}_{a}[\textbf{x}_{i}||\textbf{x}_{j}])),$ where $\textbf{W}_{a}\in\mathbb{R}^{1\times 2d_{n}}$ is a learnable parameter and $||$ denotes concatenation.
+The aggregation function then combines weighted neighbor features:
 $\textbf{h}_{i}=\textbf{W}_{h}\left(\textbf{x}_{i}+\sum_{v_{j}\in\mathcal{N}(v_{i})}\alpha_{ij}\textbf{x}_{j}\right),$ (8)
-where $\textbf{W}_{h}$ is another learnable weight matrix.
-Graph attention mechanisms have the potential to extend their application beyond nodes to edges.
-This can be achieved by substituting nodes with edges and utilizing edge features instead of node features [73].
-Furthermore, stacking of graph attention mechanisms in multiple layers allows for the utilization of output representations from one layer as input features for the subsequent layer [74].
+where $\textbf{W}_{h}$ is another learnable matrix.
+These mechanisms can also be applied to edges by treating edges as primary elements and using edge features [68].
+Stacking multiple layers of graph attention allows each layer’s output to inform subsequent layers [69].
 III-B1 Global Attention Mechanisms
-Global attention mechanisms can determine how each node calculates its attention weights across all other nodes in the graph [75].
-Global attention mechanisms can be broadly categorized into two types: quadratic attention mechanisms and linear attention mechanisms.
-Quadratic Attention Mechanisms.
-Quadratic attention mechanisms are derived from the conventional self-attention formula.
-This formula calculates attention weights by applying a softmax function to scale the dot product between the query and key vectors of each node:
-$\alpha_{ij}=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})}.$ (24)
-The computational complexity of this method is $O(N^{2})$.
-One of the pioneering studies that has introduced quadratic attention mechanisms for graph transformers is the work by Velickovic et al.[42].
-The authors proposed the use of multiple attention heads to capture different types of relations between nodes.
-Choromanski et al.[76] introduced the Graph Kernel Attention Transformer (GKAT), an approach that integrates graph kernels, structural priors, and efficient transformer architectures.
-GKAT leverages graph kernels as positional encodings to capture the structural characteristics of the graph while employing low-rank decomposition techniques to minimize the memory requirements of quadratic attention:
-$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{j})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{n})},$ (25)
-$\mathbf{z}_{i}$ $=\sum_{j=1}^{N}\alpha_{ij}(\mathbf{U}\mathbf{V}^{\top})_{ij}$
-where $\textbf{z}_{i}$ is the output vector of node $v_{i}$, $\mathbf{p}_{i}$ is the positional encoding vector of node $v_{i}$, which is computed by applying a graph kernel function to node features.
-$\mathbf{U}$ and $\mathbf{V}$ are low-rank matrices that approximate value matrix in the GKAT model.
-GKAT also introduced a novel kernel-based masking scheme to control the sparsity of the attention matrix.
+Global attention mechanisms define how each node allocates attention weights to all other nodes in the graph [70].
+They are commonly divided into two main categories: quadratic attention and linear attention.
+Quadratic Attention. Quadratic attention follows the standard self-attention formulation, using a softmax function on the scaled dot product of query and key vectors:
+$\alpha_{ij}=\frac{\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}\bigr)},$ (9)
+with a computational complexity of $O(N^{2})$.
+Veličković et al.
+introduced this approach for graph transformers with multi-head attention [37].
+Choromanski et al.
+proposed the Graph Kernel Attention Transformer (GKAT), which combines graph kernels, structural priors, and efficient attention:
+$\alpha_{ij}$ $=\frac{\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{j}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{n}\bigr)},$ (10)
+$\mathbf{z}_{i}$ $=\sum_{j=1}^{N}\alpha_{ij}(\mathbf{U}\mathbf{V}^{\top})_{ij},$
+where $\mathbf{p}_{i}$ is a positional encoding based on graph kernels, and $\mathbf{U},\mathbf{V}$ are low-rank matrices approximating the value matrix [71].
 Yun et al.
-[36] introduced Graph Transformer Networks (GTN), an approach that uses multiple layers of quadratic attention to acquire hierarchical representations of graphs.
-GTN further enhances attention computation by incorporating edge features through the use of bilinear transformations instead of concatenation.
-This innovative technique demonstrates the potential for improved graph representation learning in the field of graph analysis.
-Quadratic attention mechanisms possess a significant advantage in capturing extensive dependencies and global information within graphs, thereby offering potential benefits for various graph learning tasks.
-Nonetheless, these mechanisms have some limitations, including their computationally expensive nature, high memory consumption, and susceptibility to noise and outliers.
-Additionally, quadratic attention mechanisms may not effectively preserve and leverage local information within graphs which holds significance in certain domains and tasks.
-Linear Attention Mechanisms.
-Linear attention mechanisms employ different approximations and hashing techniques to decrease the computational complexity of self-attention from $O(N^{2})$ to $O(N)$ [77].
-These mechanisms can be categorized into two subtypes: kernel-based linear attention mechanisms and locality-sensitive linear attention mechanisms.
-Kernel-based linear attention mechanisms leverage the concept of employing kernel functions to map query and key vectors into a feature space that allows efficient computation of their inner products [78, 79].
-In the work by Katharopoulos et al.
-[80], Linear Transformer Networks (LTN) were introduced.
-LTN utilizes random Fourier features as kernel functions to approximate self-attention:
-$\alpha_{ij}$ $=\frac{\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{j})}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{n})}{\sqrt{d_{k}}})},$ (26)
-$\phi(\mathbf{x})$ $=\sqrt{\frac{2}{m}}[\cos(\omega_{1}^{\top}\mathbf{x}+b_{1}),...,\cos(\omega_{m}^{\top}\mathbf{x}+b_{m})]^{\top}.$
-Here, $\phi$ is a random Fourier feature function that maps input vector into a $m$-dimensional feature space.
-$\omega_{i}$ and $b_{i}$ are randomly sampled from a Gaussian distribution and a uniform distribution, respectively.
-Locality-sensitive linear attention mechanisms are based on the concept of employing hashing functions to partition the query and key vectors into discrete buckets that facilitate local computation of their inner products [81, 82, 83].
-Kitaev et al.
-[84] introduced the Reformer model which uses locality-sensitive hashing (LSH) as a hashing function to approximate self-attention:
-$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n\in B_{i}}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})},$ (27)
-$B_{i}$ =$=\{j|h(\mathbf{q}_{i})=h(\mathbf{k}_{j})\},$
-where $h$ is an LSH function that maps the input vector into a bucket index and $B_{i}$ is a set of nodes that share the same bucket index with node $v_{i}$.
-RGN also incorporated edge features into the attention computation using a shared query-key projection matrix.
-Linear attention mechanisms offer a significant advantage in terms of reducing computational cost and memory usage in self-attention, thereby enabling graph transformers to efficiently process large-scale graphs.
-However, they do come with certain limitations, including approximation error, hashing collision and loss of global information.
-Additionally, linear attention mechanisms may not be capable of capturing intricate or nonlinear relationships between nodes.
+introduced Graph Transformer Networks (GTN), which compute hierarchical representations using multiple layers of quadratic attention [31].
+Despite their ability to capture global relationships, these mechanisms can be memory-intensive and prone to noise, and they may overlook some local information.
+Linear Attention.
+Linear attention mechanisms lower self-attention’s complexity from $O(N^{2})$ to $O(N)$ by using approximations or hashing [72].
+They typically fall into two subtypes: kernel-based and locality-sensitive.
+Kernel-based methods map query-key vectors into a feature space for efficient dot-product computation [73, 74].
+For instance, Katharopoulos et al.
+introduced Linear Transformer Networks (LTN) with random Fourier features [75]:
+$\alpha_{ij}$ $=\frac{\exp\bigl(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{j})}{\sqrt{d_{k}}}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{n})}{\sqrt{d_{k}}}\bigr)},$ (11)
+$\phi(\mathbf{x})$ $=\sqrt{\frac{2}{m}}\bigl[\cos(\omega_{1}^{\top}\mathbf{x}+b_{1}),\dots,\cos(\omega_{m}^{\top}\mathbf{x}+b_{m})\bigr]^{\top}.$
+Locality-sensitive linear attention uses hashing to group queries and keys into buckets for local dot-product computation [76, 77, 78].
+Kitaev et al.’s Reformer model approximates self-attention with locality-sensitive hashing [79]:
+$\alpha_{ij}$ $=\frac{\exp\bigl(\tfrac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}\bigr)}{\sum_{n\in B_{i}}\exp\bigl(\tfrac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}\bigr)},$ (12)
+$B_{i}$ $=\{j\mid h(\mathbf{q}_{i})=h(\mathbf{k}_{j})\},$
+where $h$ is a locality-sensitive hash function.
+Linear attention mechanisms are more scalable for large graphs but can introduce approximation errors, hashing collisions, and partial loss of global information, potentially limiting their ability to capture complex node relationships.
 III-B2 Local Attention Mechanisms
 Local attention mechanisms determine how each node computes its attention weights over a subset of nodes in the graph [80].
-Local attention mechanisms can be broadly categorized into two types: message-passing attention mechanisms and spectral attention mechanisms.
-Message-passing Attention Mechanisms.
-Message-passing attention mechanisms are built on the foundational framework of message-passing neural networks (MPNNs) [86], which iteratively aggregate messages from neighboring nodes to compute node representations.
-To enhance MPNNs, message-passing attention mechanisms employ self-attention to compute messages or aggregation weights [61, 87, 88, 39].
-The computational complexity of this method is $O(E)$, where $E$ represents the number of edges in the graph.
-One of the pioneering works in introducing message-passing attention mechanisms for graph transformers is GraphSAGE [89].
-GraphSAGE proposes the use of mean pooling as an aggregation function and self-attention as a combination function.
-Additionally, GraphSAGE incorporates node sampling techniques to efficiently handle large-scale graphs.
-The function can be mathematically defined as:
+These mechanisms can be broadly categorized into two types: message-passing attention mechanisms and spectral attention mechanisms.
+Message-passing Attention.
+Message-passing attention mechanisms build on message-passing neural networks (MPNNs) [81], which iteratively aggregate messages from neighboring nodes to compute node representations.
+These mechanisms enhance MPNNs by using self-attention to compute messages or aggregation weights [56, 82, 83, 34].
+The computational complexity is $O(E)$, where $E$ is the number of edges.
+GraphSAGE [84] introduced message-passing attention mechanisms for graph transformers, using mean pooling as an aggregation function and self-attention as a combination function, along with node sampling for large-scale graphs.
+The function is defined as:
 $\mathbf{h}_{i}^{(l+1)}$ $=\text{ReLU}(\mathbf{W}^{(l)}[\mathbf{h}_{i}^{(l)}\|\text{MEAN}(\{\mathbf{h}_{j}^{(l)},\forall v_{j}\in\mathcal{N}(v_{i})\})])$ (13)
 $\mathbf{z}_{i}$ $=\sum_{l=0}^{L}\alpha_{il}\mathbf{h}_{i}^{(l)}.$
-[…]
-Spectral Attention Mechanisms.
-Spectral attention mechanisms are founded on the concept of transforming node features into a spectral domain, where the graph structure is encoded by eigenvalues and eigenvectors of the graph Laplacian matrix [91].
-Spectral attention mechanisms employ self-attention to calculate spectral coefficients and spectral filters [92].
-The computational complexity of this approach is $O(N)$.
+Here, $\mathbf{h}_{i}^{(l)}$ is the hidden vector of node $v_{i}$ at layer $l$, $\mathbf{W}^{(l)}$ is a linear transformation matrix, $\|$ is concatenation, and MEAN is mean pooling.
+$\alpha_{il}$ is the attention weight between node $v_{i}$ and layer $l$.
+Javaloy et al.
+[85] proposed L-CAT, using self-attention for both aggregation and combination, and incorporating edge features with bilinear transformations.
+Message-passing attention mechanisms effectively preserve and leverage local graph information but face scalability issues, including limited expressiveness and dependence on graph connectivity.
+Their ability to capture long-range dependencies and global graph information is also limited.
+Spectral Attention.
+Spectral attention mechanisms transform node features into a spectral domain, encoding graph structure with eigenvalues and eigenvectors of the graph Laplacian matrix [86].
+They use self-attention to calculate spectral coefficients and filters [87], with a computational complexity of $O(N)$.
 Wang et al.
-[22] proposed graph isomorphism networks (GIN) as a spectral attention mechanism for graph transformers, employing an expressive aggregation function of sum pooling and self-attention.
-GIN also integrates a unique graph readout scheme via set functions to encapsulate global graph characteristics.
-The equations of GIN are as follows:
+[22] proposed Graph Isomorphism Networks (GIN) as a spectral attention mechanism for graph transformers, utilizing sum pooling and self-attention.
+GIN integrates a graph readout scheme via set functions to encapsulate global graph characteristics, defined as:
 $\mathbf{h}_{i}^{(l+1)}$ $=\text{MLP}^{(l)}((1+\epsilon^{(l)})\mathbf{h}_{i}^{(l)}+\sum_{j\in\mathcal{N}(v_{i})}\mathbf{h}_{j}^{(l)}),$ (14)
@@
 $\mathbf{z}_{G}$ $=\text{READOUT}(\{\mathbf{z}_{i},\forall v_{i}\in V\}).$
-Here, $\text{MLP}^{(l)}$ is a multi-layer perceptron, $\epsilon^{(l)}$ is a learnable parameter, READOUT indicates a set function that aggregates node output vectors into a graph output vector.
-$\mathbf{z}_{G}$ is the output vector of graph $G$.
-In addition, Nguyen et al.
-[93] introduced UGformer, a self-attention-based method for the spectral coefficient computation of each node using eigenvalues and eigenvectors of the Laplacian graph matrix.
-UGformer further integrates edge features into spectral computation via bilinear transformations.
-Spectral attention mechanisms possess the ability to incorporate the structural information of a graph into the spectral domain, thereby offering potential benefits for certain tasks or domains.
-However, they are also accompanied by certain limitations, including high computational cost, memory consumption and sensitivity to graph size and density.
+Here, $\text{MLP}^{(l)}$ is a multi-layer perceptron, $\epsilon^{(l)}$ is a learnable parameter, and READOUT aggregates node output vectors into a graph output vector $\mathbf{z}_{G}$.
+Nguyen et al.
+[88] introduced UGformer, which uses self-attention for spectral coefficient computation and incorporates edge features via bilinear transformations.
+Spectral attention mechanisms can incorporate graph structural information into the spectral domain, offering benefits for certain tasks.
+However, they face limitations such as high computational cost, memory consumption, and sensitivity to graph size and density.
 IV Taxonomy of Graph Transformers
-The past few years have witnessed a surge of interest in graph transformers.
-This section dives into four key categories dominating the current literature: shallow, deep, scalable, and pre-trained graph transformers.
-By analyzing representative models within each category, we aim to establish valuable guidelines for designing effective graph transformers.
+Recent years have seen a surge of interest in graph transformers.
+This section examines four key categories in the literature—shallow, deep, scalable, and pre-trained graph transformers—organized to highlight distinctions between these categories.
+For example, shallow and deep graph transformers are compared based on model depth and ability to capture dependencies, while scalable and pre-trained transformers focus on handling large-scale graphs and leveraging unlabeled data, respectively.
+By expanding on specific examples of Graph Transformer models in each category, we aim to elucidate the architectural choices, strengths, and limitations that guide practitioners in selecting suitable approaches for their tasks.
+By analyzing representative models in each category, we aim to provide guidelines for designing effective graph transformers.
 IV-A Shallow Graph Transformers
-Shallow graph transformers represent a class of GNNs that leverage the power of self-attention to acquire node representations from data structured in graphs.
-Inspired by transformer models, which effectively capture long-range dependencies in sequential data through self-attention, shallow graph transformers extend this concept to graph data by computing self-attention weights based on both node features and graph topology [94].
-The primary objective of shallow graph transformers is to achieve exceptional performance while minimizing computational complexity and memory usage.
-Shallow graph transformers can be seen as a generalization of graph attention networks (GAT) [42].
-GAT use a multi-head attention mechanism to calculate node embeddings.
-However, GAT has some limitations, such as the inability to model edge features and lack of diversity among attention heads [95].
-Several GAT extensions have been proposed in the literature to address these issues.
-For example, GTN by Yun et al.
+Shallow graph transformers leverage self-attention to derive node representations from graph-structured data.
+Inspired by transformer models that capture long-range dependencies in sequential data, these transformers extend this concept to graph data by computing self-attention weights based on both node features and graph topology [89].
+The primary goal of shallow graph transformers is to achieve high performance while minimizing computational complexity and memory usage.
+Unlike deep transformers, these models are optimized for efficiency and typically involve fewer layers, focusing on lightweight computation for smaller graphs or real-time applications.
+These transformers can be viewed as a generalization of graph attention networks (GAT) [37], which use multi-head attention to compute node embeddings.
+However, GATs have limitations, such as the inability to model edge features and a lack of diversity among attention heads [90].
+To address these issues, several extensions have been proposed.
+For instance, GTN by Yun et al.
 [31] introduces edge-wise self-attention to incorporate edge information into node embeddings.
-Ahmad et al.[63] proposed the graph attention transformer encoder (GATE), which applies a masked self-attention mechanism to learn different attention patterns for different nodes.
-GATE also uses a position-wise feed-forward network and dropout to enhance model capacity and generalization.
+Similarly, Ahmad et al.
+[58] proposed the Graph Attention Transformer Encoder (GATE), which applies a masked self-attention mechanism to learn distinct attention patterns for different nodes.
+Such innovations demonstrate the adaptability of shallow graph transformers, but their limited depth constrains their ability to capture complex graph dependencies, which are better addressed by deeper models.
+In practice, shallow models often excel in lower-resource or time-sensitive scenarios due to their reduced computational overhead and simpler optimization.
+However, their smaller capacity may limit performance on highly complex tasks, making them most effective when graph structures are relatively simple or when inference speed is a priority.
 The summary of shallow graph transformer methods is given in Table II.
-Shallow graph transformers are efficient and adaptable capable of handling various graph learning tasks and different types of graphs, but their lack of depth and recurrence may limit their ability to capture complex dependencies.
-Their performance can also be influenced by the choice of mask matrix and the number of attention heads, indicating a need for further research on their optimal design and regularization.
 TABLE II: Shallow graph transformers
-Model Graph inductive bias Graph attention mechanisms Application Dataset Code availability
+Model Graph inductive bias Graph attention mechanisms Application Datasets Code availability
 Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
@@
 GATE [58] No No No Yes (linear) No Cross-lingual relation and event extraction ACE 2005 corpus Yes
-PiT [96] No No No Yes (quadratic) No Image classification, object detection ImageNet, COCO, ImageNet-A Yes
-WDPruning [97] No No No Yes (quadratic) No Image classification and image retrieval CIFAR-10 and ILSVRC-12 No
-SVT [98] No No No Yes (quadratic) No Image classification Synthetic data and CIFAR-10 No
 Gubelmann et al. [91] No No No Yes No Natural language inference (NLI) SNLI and MNLI No
 TADDY [92] Yes (global and local) No No Yes (linear) No Anomaly detection in dynamic graphs UCI Messages, Enron, DBLP, Facebook, Twitter and Reddit. No
-Tree Transformer [38] Yes (local) No No Yes (quadratic) No Program translation and semantic parsing For2Lam, CoffeeScript-JavaScript, JOBS, GEO, ATIS Yes
+Tree Transformer [33] Yes (local) No No Yes (linear) No Program translation and semantic parsing For2Lam, CoffeeScript-JavaScript, JOBS, GEO, ATIS Yes
 GAT-POS  [38] Yes (local) No No Yes (linear) Yes (message passing) Node classification Chameleon, Squirrel, Actor, Cora, Citeseer, Pubmed No
@@
 IV-B Deep Graph Transformers
-Deep graph transformers consist of multiple self-attention layers stacked on top of each other, with optional skip connections, residual connections or dense connections between layers [102].
-They are designed to achieve higher performance with increased model depth and complexity [103].
-Deep graph transformers extend shallow graph transformers by applying self-attention layers to node features and graph topology hierarchically.
-However, deep graph transformers also face several challenges that need to be addressed.
-One challenge is the difficulty of training deeper models, which can be mitigated by employing techniques, such as PairNorm introduced in DeeperGCN [104].
-Another challenge is the over-smoothing problem, which can be addressed by using a gated residual connection and a generalized convolution operator as proposed in DeeperGCN.
-Additionally, the disappearance of global attention capacity and the lack of diversity among attention heads are challenges that can be tackled by approaches like DeepGraph [94].
-DeepGraph incorporates substructure tokens and local attention to improve the focus and diversity of global attention.
-Deep graph transformers, while complex, can achieve top-tier results on various graph learning tasks and adapt to different types of graphs and domains.
-However, their high computational cost, difficulty in optimization, and sensitivity to hyperparameters necessitate further research for optimal design and training.
+Deep graph transformers consist of multiple stacked self-attention layers with optional skip, residual, or dense connections between layers [94].
+They aim for higher performance through increased model depth and complexity [95, 96], applying self-attention layers hierarchically to node features and graph topology.
+Unlike shallow models, these transformers can capture complex, multi-level dependencies in large and heterogeneous graphs, making them suitable for tasks requiring deep feature extraction.
+Training deeper models is challenging, often requiring techniques like PairNorm from DeeperGCN to mitigate difficulties [97].
+Over-smoothing can be addressed using gated residual connections and generalized convolution operators, also proposed in DeeperGCN.
+Other challenges, such as the loss of global attention capacity and lack of diversity among attention heads, can be tackled by approaches like DeepGraph [89], which uses substructure tokens and local attention to enhance focus and diversity.
+Although deep graph transformers achieve state-of-the-art performance on many graph tasks, their optimization remains sensitive to hyperparameter tuning, computational costs, and architectural choices, which require further study.
+Deeper architectures generally exhibit stronger representational power, enabling them to model complex graph relationships and long-range dependencies.
+Nonetheless, this comes at a higher computational cost, greater risk of vanishing gradients, and potential overfitting.
+Successful deep transformer designs often incorporate specialized normalization and regularization layers to preserve training stability and manage computational overhead.
 Table III summarizes the methods of deep graph transformers.
@@
 Node Positional encoding Edge Structural Encoding Message-Passing Bias Global attention Local Attention
-HG-Transformer [102] No No No Yes (linear) No Multi-label text classification RCV1, RCV1-2K, AmazonCat-13K No
-Unfolded Transformer [103] No No No Yes (quadratic) No Graph neural networks and NLP IMDB and SST2 Yes
 DeepGraph [89] Yes (relative) No No Yes (linear) Yes (message passing) Graph representation learning PCQM4M-LSC, ZINC, CLUSTER, PATTERN Yes
@@
 GraphiT [41] Yes (local and global) No No Yes (quadratic) Yes (message passing) Graph representation learning for classification and regression tasks MUTAG, PROTEINS, PTC, NCI1, ZINC Yes
-Tree Transformer [48] Yes (global and local) No No Yes (quadratic) Yes (message passing) Code representation learning Python150k and JavaScript150k Yes
-MagLapNet [49] Yes (local and global) No No Yes (quadratic) Yes (spectral) Graph learning tasks such as correctness testing of sorting networks and source code understanding Synthetic graphs, OGB Code2 Yes
-SAT [52] Yes (local and global) No No Yes (quadratic) Yes (message passing) Graph representation learning for graph and node property prediction ZINC, CLUSTER, PATTERN, OGBG-PPA, OGBG-CODE2 Yes
-Edge Transformer [107] Yes Yes No Yes (quadratic) No Improve systematic generalization in natural language understanding and relational reasoning tasks CLUTRR, CFQ and COGS Yes
+MagLapNet [44] Yes (local and global) No No Yes (quadratic) Yes (spectral) Source code understanding Synthetic graphs, OGB Code2 Yes
+SAT [47] Yes (local and global) No No Yes (quadratic) Yes (message passing) Node property prediction ZINC, CLUSTER, PATTERN, OGBG-PPA, OGBG-CODE2 Yes
+Edge Transformer [100] Yes Yes No Yes (quadratic) No Generalization in natural language understanding and relational reasoning tasks CLUTRR, CFQ and COGS Yes
 Cai et al. [57] Yes Yes No Yes (quadratic) No Graph-to-sequence learning LDC2015E86, LDC2017T10, WMT16 English-German, WMT16 English-Czech Yes
 SAN [61] Yes (local and global) No No Yes (quadratic) Yes (spectral) Graph representation learning ZINC, PATTERN, CLUSTER, MolHIV, MolPCBA Yes
-KMCA [79] No No No Yes (quadratic) Yes (message passing) Time series analysis Synthetic, FordA, uWave, IMDB, ECG No
-ABT-MPNN [87] No No Yes (pre-processing) Yes (quadratic and linear) No Molecular property prediction and biological activity prediction Johnson et al.
-[108], Tox21, ClinTox, ToxCast, HIV, QM8, ESOL, FreeSolv, Lipophilicity and QM8.
-Yes
-UGformer [93] No No No Yes (quadratic) No Graph classification and text classification COLLAB, IMDB-B, IMDB-M, DD, PROTEINS, MUTAG, PTC, MR, R8, R52, Ohsumed Yes
+ABT-MPNN [82] No No Yes (pre-processing) Yes (quadratic and linear) No Molecular property prediction Johnson et al. [101], Tox21, ClinTox, ToxCast, HIV, QM8, ESOL, FreeSolv, Lipophilicity and QM8. Yes
+UGformer [88] No No No Yes (quadratic) No Graph classification COLLAB, IMDB-B, IMDB-M, DD, PROTEINS, MUTAG, PTC, MR, R8, R52, Ohsumed Yes
 DTI-GTN [102] No Yes Yes (interleaving) Yes (quadratic) No Drug-target interaction prediction Peng et al. [103] dataset No
@@
 IV-C Scalable Graph Transformers
-Scalable graph transformers are a category of graph transformers that tackle the challenges of scalability and efficiency when applying self-attention to large-scale graphs [114, 39, 53, 115].
-These transformers are specifically designed to reduce computational cost and memory usage while maintaining or improving performance.
-To achieve this, various techniques are employed to reduce the complexity of self-attention, such as sparse attention, local attention, and low-rank approximation [12, 115].
-Scalable graph transformers can be regarded as an enhancement of deep graph transformers addressing challenges, such as over-smoothing and limited capacity of global attention.
-Several scalable graph transformer models have been proposed to enhance the scalability and efficiency of graph transformers.
-For instance, Rampášek et al.
-[39] introduced GPS, use low-rank matrix approximations to reduce computational complexity, and achieve state-of-the-art results on diverse benchmarks.
-GPS decouples local real-edge aggregation from a fully-connected transformer and incorporates different positional and structural encodings to capture graph topology.
-It also offers a modular framework that supports multiple encoding types and mechanisms for local and global attention.
+Scalable graph transformers address the challenges of scalability and efficiency in applying self-attention to large-scale graphs [107, 34, 48, 108].
+These models reduce computational cost and memory usage while maintaining or improving performance by employing techniques such as sparse attention, local attention, and low-rank approximation [12, 108].
+Compared to deep transformers, scalable models are optimized for handling large datasets and real-world applications with limited resources.
+Several models have been proposed to enhance scalability and efficiency.
+Rampášek et al.
+[34] introduced GPS, which uses low-rank matrix approximations to reduce computational complexity and achieve state-of-the-art results on diverse benchmarks.
+GPS decouples local real-edge aggregation from a fully-connected transformer and incorporates various positional and structural encoding to capture graph topology, offering a modular framework for local and global attention mechanisms.
 Cong et al.
-[116] developed DyFormer, a dynamic graph transformer that utilizes substructure tokens and local attention to enhance the focus and diversity of global attention.
-DyFormer employs a temporal union graph structure and a subgraph-based node sampling strategy for efficient and scalable training.
-Scalable graph transformers are an innovative and efficient category of graph transformers that excel in handling large-scale graphs while minimizing computational cost and memory usage.
-However, scalable graph transformers face certain limitations, including the trade-off between scalability and expressiveness, the challenge of selecting optimal hyperparameters and encodings, and the absence of theoretical analysis regarding their convergence and stability.
-Consequently, further investigation is required to explore optimal designs and evaluations of scalable graph transformers for various applications.
-For a comprehensive overview of scalable graph transformer methods, please refer to Table IV.
+[109] developed DyFormer, a dynamic graph transformer that utilizes substructure tokens and local attention to enhance global attention diversity.
+While scalable graph transformers address key issues of memory and computation, they often involve trade-offs between scalability and expressiveness, requiring further research on optimal encoding and task-specific adaptations.
+Scalable models typically strike a balance between performance and resource demands, employing approximations or partitioning methods to handle large datasets without exhaustive computation.
+However, aggressive approximation can lead to reduced expressiveness, and domain-specific strategies may be needed to recover performance in certain tasks or graph structures.
+For a comprehensive overview, see Table IV.
 TABLE IV: Scalable graph transformers
@@
 NodeFormer [48] No No Yes (relational bias) Yes (linear) No Node classification Citation networks, OGB-Proteins, Amazon2M, Mini-ImageNet, 20News-Groups Yes
-DIFFORMER [115] No No No Yes (linear) Yes (spectral) Node classification, image/text classification, spatial-temporal dynamics prediction Cora, Citeseer, Pubmed, ogbn-Proteins, Pokec, CIFAR, STL, 20News Yes
+DIFFORMER [108] No No No Yes (linear) Yes (spectral) Node classification Cora, Citeseer, Pubmed, ogbn-Proteins, Pokec, CIFAR, STL, 20News Yes
 DyFormer [109] Yes (global) Yes (global) Yes (pre-processing) Yes (quadratic) No Dynamic graph representation learning RDS, UCI, Yelp, ML-10M, SNAP-Wikipedia, SNAP-Reddit No
-Graph ViT/MLP-Mixer [117] Yes No No No Yes (message passing) Graph representation learning ZINC, MNIST, CIFAR10, MolTOX21, MolHIV, Peptide-func, Peptide-struct, CSL, EXP, SR25, LRGB, TreeNeighbourMatch Yes
 DiGress [110] No Yes No Yes Yes Graph generation QM9, MOSES, GuacaMol, SBM, planar graphs Yes
-EXPHORMER [119] Yes (local and global) Yes (local and global) Yes (interleaving) Yes (linear) Yes (spectral) Graph learning and representation tasks CIFAR10, MalNet-Tiny, MNIST, CLUSTER, PATTERN, PascalVOC-SP, COCO-SP, PCQM-Contact, ogbn-arxiv, Yes
-EGT [120] Yes (global) Yes (global) No Yes (quadratic) No Node and edge classification, graph classification and regression, transfer learning.
-PATTERN, CLUSTER, TSP, MNIST, CIFAR10, ZINC, PCQM4M, PCQM4Mv2, MolPCBA and MolHIV.
-Yes
+EXPHORMER [111] Yes (local and global) Yes (local and global) Yes (interleaving) Yes (linear) Yes (spectral) Graph classification CIFAR10, MalNet-Tiny, MNIST, CLUSTER, PATTERN, PascalVOC-SP, COCO-SP, PCQM-Contact, ogbn-arxiv, Yes
+EGT [112] Yes (global) Yes (global) No Yes (quadratic) No Node and edge classification PATTERN, CLUSTER, TSP, MNIST, CIFAR10, ZINC, PCQM4M, PCQM4Mv2, MolPCBA and MolHIV. Yes
 GRIT [113] Yes No No Yes (quadratic) No Graph regression and classification ZINC, MNIST, CIFAR10, PATTERN, CLUSTER, Peptides-func, Peptides-struct, ZINC-full, PCQM4Mv2 Yes
 HSGT [114] No Yes (local, based on shortest path distance) No Yes (quadratic) Yes (message passing, based on neighborhood sampling) Node classification Cora, CiteSeer, PubMed, Amazon-Photo, ogbn-arxiv, ogbn-proteins, ogbn-products, Reddit, Flickr, Yelp No
-Higher-Order Transformers [123] No No No Yes (quadratic and linear) Yes (message passing) Graph and hypergraph modeling Synthetic chains, PCQM4M-LSC, Jets, Delaunay, Hyperedge prediction Yes
-EGAT [73] No No Yes (interleaving) Yes (linear) No Graph representation learning and molecular property prediction AMLSim, Cora, Citeseer, PubMed No
-L-CAT [90] No No No Yes (linear) No Node classification CSBM, Amazon, Facebook, GitHub, TwitchEN, Coauthor Physics, OGB (arxiv, products, mag, proteins) Yes
+EGAT [68] No No Yes (interleaving) Yes (linear) No Molecular property prediction AMLSim, Cora, Citeseer, PubMed No
 Point Transformer [83] Yes (relative) No No No Yes (message passing) 3D point cloud processing S3DIS, ModelNet40, ShapeNetPart Yes
@@
 HierGAT [115] Yes (local) No No Yes (linear) No Entity Resolution Magellan and WDC product matching datasets Yes
-Geo-ER [125] No No No Yes (linear) Yes (message passing) Geospatial Entity Resolution OSM, FSQ, Yelp Yes
-GTA [126] No No No Yes (linear) Yes (message passing) Multivariate time series anomaly detection in IoT SWaT, WADI, SMAP, MSL Yes
 GraTransDRP [116] Yes No No Yes (linear) No Drug response prediction GDSC Yes
@@
 RelTR [120] No No No Yes (linear) No Scene graph generation Visual Genome, Open Images V6, VRD Yes
-BSTG-Trans [132] Yes (local and global) Yes (local and global) No Yes (linear) Yes (message Passing) Long-term Pose Forecasting Human3.6M, HumanEva-I, Human360K Yes
+BSTG-Trans [121] Yes (local and global) Yes (local and global) No Yes (linear) Yes (message Passing) Pose forecasting Human3.6M, HumanEva-I, Human360K Yes
 GraphSum [122] Yes (local and global) No No Yes (linear) Yes (message passing) Abstractive MDS WikiSum and MultiNews Yes
-KG-Transformer [134] No No Yes (pre-processing) Yes (linear) No Text generation from knowledge graphs UMLS, PubMed, 2010 i2b2/VA, JNLPBA, BC5CDR, GAD, EU-ADR No
+KG-Transformer [123] No No Yes (pre-processing) Yes (linear) No Text generation UMLS, PubMed, 2010 i2b2/VA, JNLPBA, BC5CDR, GAD, EU-ADR No
 IV-D Pre-trained Graph Transformers
-Pre-trained graph transformers utilize large-scale unlabeled graphs to acquire transferable node embeddings [135].
-These embeddings can be fine-tuned for downstream tasks with scarce labeled data that address the challenges of data scarcity and domain adaptation in graph learning tasks [136, 137].
-These transformers are similar to pre-trained large language models (LLMs) and are trained on graph datasets using self-supervised learning objectives, such as masked node prediction [138], edge reconstruction [139], and graph contrastive learning [140].
-These objectives aim to encapsulate the inherent properties of graph data independently of external labels or supervision [141].
-The pre-trained model can be fine-tuned on a specific downstream task with a smaller or domain-specific graph dataset by incorporating a task-specific layer or loss function and optimizing it on labeled data [142].
-This allows the pre-trained model to transfer the knowledge acquired from the large-scale graph dataset to the subsequent task, giving better performance compared to the training from scratch [142].
-Pre-trained graph transformers face some challenges, such as the selection of appropriate pre-training tasks, domain knowledge incorporation, heterogeneous information integration, and pre-training quality evaluation [143].
-To address these issues, KPGT [144] and KGTransformer [145] have been proposed.
-KPGT leverages additional domain knowledge for pre-training, while KGTransformer serves as a uniform Knowledge Representation and Fusion (KRF) module in diverse tasks.
-Despite their power and flexibility, pre-trained graph transformers encounter issues related to graph data heterogeneity and sparsity, domain adaptation, model generalization and performance interpretation.
-A summary of pre-trained graph transformer methods is provided in Table V.
+Pre-trained graph transformers leverage large-scale unlabeled graphs to acquire transferable node embeddings, which can be fine-tuned for downstream tasks with limited labeled data [124, 125, 126].
+These models are analogous to pre-trained language models, enabling knowledge transfer and reducing data dependence by using general-purpose embeddings.
+Similar to pre-trained large language models (LLMs), these transformers are trained using self-supervised learning objectives like masked node prediction [127], edge reconstruction [128], and graph contrastive learning [129], capturing the inherent properties of graph data without external labels [130].
+The pre-trained model can then be fine-tuned on specific tasks using a smaller or domain-specific graph dataset, incorporating task-specific layers or loss functions, leading to better performance compared to training from scratch.
+Challenges for pre-trained graph transformers include selecting appropriate pre-training tasks, incorporating domain knowledge, integrating heterogeneous information, and evaluating pre-training quality [131].
+Models like KPGT [132], which uses additional domain knowledge, and KGTransformer [133], which acts as a uniform Knowledge Representation and Fusion (KRF) module, have been proposed to address these issues.
+While pre-trained models are versatile, their application is constrained by challenges related to domain-specific adaptations, evaluation consistency, and interpretability of embeddings.
+Further work is needed to standardize benchmarks and improve transfer learning capabilities.
+Compared to shallow or scalable transformers, pre-trained models often excel at adapting to new tasks with minimal labeled data, leveraging rich representations gleaned from large-scale training.
+However, this advantage can be limited if the pre-training domain diverges significantly from the target domain or if specialized graph structures require model re-parameterization.
+For a comprehensive overview, see Table V.
 TABLE V: Pre-trained graph transformers
@@
 HEAT [128] Yes (local) No No Yes (linear) Yes (message passing) Structured reconstruction SpaceNet Challenge and Structured3D Yes
-FormNetV2 [140] No No No Yes (linear) Yes (message passing) Form document information extraction FUNSD, CORD, SROIE No
 MPG [131] No No No Yes (linear) Yes (spectral) Drug discovery MoleculeNet, DDI, DTI Yes
@@
 CoVGT [135] No No No Yes (linear) Yes (message-passing) Video Question Answering NExT-QA, TGIF-QA, TGIF-QA-R, STAR-QA, Causal-VidQA, MSRVTT-QA Yes
-G-Adapter [148] No Yes (local) No Yes (quadratic) No Molecular graph tasks MolHIV, MolPCBA, FreeSolv, ESOL, BBBP, Estrogen-$\alpha$, Estrogen-$\beta$, MetStablow, MetStabhigh Yes
+G-Adapter [136] No Yes (local) No Yes (quadratic) No Molecular graph tasks MolHIV, MolPCBA, FreeSolv, ESOL, BBBP, $\alpha, Estrogen$-$\beta$, MetStablow, MetStabhigh Yes
 kgTransformer [137] No No Yes (interleaving) Yes (quadratic) No Knowledge graph reasoning for complex logical queries FB15k-237, NELL995 Yes
-Li et al. [149] Yes (local and global) No No Yes (linear) Yes (message passing) Movement synchrony estimation in autism therapy interventions PT13 and Human3.6MS No
-Pellegrini et al. [150] No Yes No Yes (linear) No Disease prediction using graph convolutional networks TADPOLE and MIMIC-III Yes
-Video Graph Transformer (VGT) [151] No No No Yes (linear) Yes (message passing) Video Question Answering (VideoQA) NExT-QA, TGIF-QA, MSRVTT-QA Yes
-Tan et al. [152] No No No Yes (linear) No Few-shot node classification CoraFull, ogbn-arxiv, CiteSeer, Cora No
-GPT-GNN [60] No Yes Yes No Yes Graph representation learning and graph mining OAG, Amazon, Reddit, OAG (citation) Yes
-IV-E Design Guide for Effective Graph Transformers
-Developing effective graph transformers requires meticulous attention to detail and careful consideration.
-This guide provides general principles and tips for designing graph transformers for various scenarios and tasks.
-  • Choose the appropriate type of graph transformers based on the nature and complexity of your graph data and tasks.
-For simple and small graph data, a shallow graph transformer with a few layers may suffice.
-For complex and large graph data, a deep graph transformer with many layers can learn more expressive representations.
-For dynamic or streaming graph data, a scalable graph transformer is more efficient.
-Pre-trained graph transformers are more suitable for sparse or noisy graph data.
-  • Design suitable structural and positional encodings for your graph data.
-These encodings capture the structure of graphs and are added to input node or edge features before feeding them to transformer layers.
-The choice of encodings depends on the characteristics of the graph data, such as directionality, weight, and homogeneity.
-The careful design of these encodings ensures their informativeness.
-  • Optimize the self-attention mechanism for your graph data.
-Self-attention mechanisms compute attention scores among all pairs of nodes or edges in the graph, capturing long-range dependencies and interactions.
-However, it introduces challenges like computational complexity, memory consumption, overfitting, over-smoothing, and over-squashing.
-Techniques like sampling, sparsification, partitioning, hashing, masking, regularization, and normalization can be employed to address these challenges and improve the quality and efficiency of the self-attention mechanism.
-  • Utilize pre-training techniques to enhance the performance of graph transformers.
-Pre-training techniques leverage pre-trained models or data from other domains or tasks transferring knowledge or parameters to a specific graph learning task.
-Methods like fine-tuning, distillation, and adaptation can be used to adapt pre-trained models or data.
-Utilizing pre-training techniques is particularly beneficial when a large amount of pre-training data or resources are available.
+Li et al. [137] Yes (local and global) No No Yes (linear) Yes (message passing) Movement synchrony estimation PT13 and Human3.6MS No
+Pellegrini et al. [138] No Yes No Yes (linear) No Disease prediction TADPOLE and MIMIC-III Yes
+Video Graph Transformer (VGT) [139] No No No Yes (linear) Yes (message passing) Video Question Answering NExT-QA, TGIF-QA, MSRVTT-QA Yes
+Tan et al. [140] No No No Yes (linear) No node classification CoraFull, ogbn-arxiv, CiteSeer, Cora No
+GPT-GNN [55] No Yes Yes No Yes Graph representation learning OAG, Amazon, Reddit, OAG (citation) Yes
 V Application Perspectives of Graph Transformers
-Graph transformers are finding applications in various domains that involve interconnected data.
-This section delves into their applications for graph-related tasks, categorized by the level of analysis: node-level, edge-level and graph-level.
-Beyond these core tasks, graph transformers are also making strides in applications that handle text, images, and videos, where data can be effectively represented as graphs for analysis.
+Graph transformers are increasingly utilized in domains involving interconnected data.
+This section explores their applications in graph-related tasks, categorized by analysis level: node-level, edge-level, and graph-level.
+Additionally, graph transformers are advancing in text, image, and video applications, where data is represented as graphs for enhanced analysis.
 V-A Node-level Tasks
@@
 V-A1 Protein Structure Prediction
-In the field of bioinformatics, graph transformers have demonstrated substantial potential in Protein Structure Prediction (PSP) [154].
+In bioinformatics, graph transformers have shown significant potential in Protein Structure Prediction (PSP) [142].
 Gu et al.
-[155] introduced HEAL, which employs hierarchical graph transformers on super-nodes that imitate functional motifs to interact with nodes in the protein graph, effectively capturing structural semantics.
+[143] introduced HEAL, which utilizes hierarchical graph transformers on super-nodes that mimic functional motifs to interact with nodes in the protein graph, capturing structural semantics.
 Pepe et al.
-[156] used Geometric Algebra (GA) modelling to introduce a new metric based on the relative orientations of amino acid residues, serving as an additional input feature to a graph transformer, assisting in the prediction of the 3D coordinates of a protein.
+[144] employed Geometric Algebra (GA) modeling to introduce a metric based on the relative orientations of amino acid residues, serving as an input feature to a graph transformer for predicting protein 3D coordinates.
 Chen et al.
-[157] proposed gated-graph transformers integrating node and edge gates within a graph transformer framework to regulate information flow during graph message passing, proving beneficial in predicting the quality of 3D protein complex structures.
-Despite the encouraging outcomes, various challenges persist, such as the complexity of protein structures, scarcity of high-quality training data, and substantial computational resource [158].
-Further investigation is necessary to address these challenges and enhance the precision of these models.
+[145] proposed gated-graph transformers that integrate node and edge gates to regulate information flow during graph message-passing, aiding in predicting the quality of 3D protein complex structures.
+Despite promising results, challenges such as protein structure complexity, limited high-quality training data, and high computational demands persist [146].
+Further research is needed to address these challenges and improve model accuracy.
 V-A2 Entity Resolution
-Entity Resolution (ER) is a crucial task in data management that aims to identify and link disparate representations of real-world entities from diverse sources [159].
-Recent research has highlighted the efficacy of graph transformers in ER.
-For example, Yao et al.
-[160] proposed Hierarchical Graph Attention Networks (HierGAT) integrating the self-attention mechanism and graph attention network mechanism to capture and leverage the relationships between different ER decisions, leading to substantial enhancements over conventional approaches.
+Entity Resolution (ER) is essential in data management, aiming to identify and link disparate representations of real-world entities from various sources [147].
+Recent studies have shown the effectiveness of graph transformers in ER.
+Yao et al.
+[148] introduced Hierarchical Graph Attention Networks (HierGAT), integrating self-attention and graph attention mechanisms to leverage relationships between different ER decisions, significantly improving over traditional methods.
 Ying et al.
-[161] extended the standard transformer architecture and introduced several straightforward yet powerful structural encoding techniques to enhance the modeling of graph-structured data.
-Despite facing challenges related to data complexity and structural information encoding, these techniques have exhibited promising outcomes in terms of enhanced performance, scalability, and accuracy.
-Furthermore, Dou et al.[162] proposed the Hybrid Matching Knowledge for Entity Matching (GTA) method improves the transformer for representing relational data by integrating additional hybrid matching knowledge acquired through graph contrastive learning on a specially designed hybrid matching graph.
-This approach has also demonstrated promising results by effectively boosting the transformer for representing relational data and surpassing existing entity matching frameworks.
+[149] extended the transformer architecture with structural encoding techniques to better model graph-structured data, demonstrating enhanced performance, scalability, and accuracy.
+Dou et al.
+[150] proposed the Hybrid Matching Knowledge for Entity Matching (GTA) method, which improves transformers for relational data by integrating hybrid matching knowledge through graph contrastive learning.
+This approach has shown promising results, surpassing existing entity matching frameworks.
+Despite challenges related to data complexity and structural information encoding, these methods have proven effective.
 V-A3 Anomaly Detection
-Graph transformers are valuable tools for anomaly detection, especially in dynamic graphs and time series data [65, 163].
-They tackle key challenges like encoding information for unattributed nodes and extracting discriminative knowledge from spatial-temporal dynamic graphs.
-Liu et al [164] proposed TADDY, a transformer-based Anomaly Detection framework, enhancing node encoding to represent each node’s structural and temporal roles in evolving graph streams.
-Similarly, Xu et al.
-[165] proposed the Anomaly Transformer which uses an Anomaly-Attention mechanism to measure association discrepancy and employs a minimax strategy to enhance normal-abnormal differentiation.
+Graph transformers are crucial for anomaly detection, particularly in dynamic graphs and time series data [60, 151].
+They address challenges such as encoding information for unattributed nodes and extracting discriminative knowledge from spatial-temporal dynamic graphs.
+Liu et al.
+[152] introduced TADDY, a transformer-based anomaly detection framework that enhances node encoding to represent structural and temporal roles in evolving graph streams.
+Xu et al.
+[153] proposed the Anomaly Transformer, which uses an Anomaly-Attention mechanism to measure association discrepancy and a minimax strategy to differentiate normal and abnormal data.
 Chen et al.
-[166] proposed the GTA framework for multivariate time series anomaly detection incorporates graph structure learning, graph convolution, and temporal dependency modeling with a transformer-based architecture.
+[154] developed the GTA framework for multivariate time series anomaly detection, incorporating graph structure learning, graph convolution, and temporal dependency modeling with a transformer-based architecture.
 Tuli et al.
-[167] developed TranAD, a deep transformer network for anomaly detection in multivariate time series data, showing efficient anomaly detection and diagnosis in modern industrial applications.
-Despite their effectiveness, further research is needed to enhance their performance and applicability across different domains.
+[155] created TranAD, a deep transformer network for anomaly detection in multivariate time series, demonstrating efficient anomaly detection and diagnosis in industrial applications.
+Although effective, further research is required to enhance performance and applicability across different domains [156].
 V-B Edge-level Tasks
@@
 V-B1 Drug-Drug Interaction Prediction
-Graph transformers have been increasingly employed in the prediction of Drug-Drug Interactions (DDIs) owing to their capability to adeptly model the intricate relationships between drugs and targets [170].
+Graph transformers have been increasingly utilized for predicting Drug-Drug Interactions (DDIs) due to their ability to model complex drug-target relationships [159].
 Wang et al.
-[109] proposed a method which uses a line graph with drug-protein pairs as vertices and a graph transformer network (DTI-GTN) for the purpose of forecasting drug-target interactions.
+[102] introduced a method using a line graph with drug-protein pairs as vertices and a graph transformer network (DTI-GTN) to predict drug-target interactions.
 Djeddi et al.
-[171] proposed a novel approach named DTIOG for the prediction of DTIs, leveraging a Knowledge Graph Embedding (KGE) strategy and integrating contextual information derived from protein sequences.
-Despite the encouraging outcomes, these approaches encounter challenges such as overlooking certain facets of the intermolecular information and identifying potential interactions for newly discovered drugs [172].
-Nevertheless, findings from multiple studies indicate that graph transformers can proficiently anticipate DDIs and surpass the performance of existing algorithms [173].
+[160] developed DTIOG, a novel approach leveraging Knowledge Graph Embedding (KGE) and contextual information from protein sequences to predict DTIs.
+Despite promising results, challenges remain, such as neglecting certain intermolecular information and identifying interactions for new drugs [161].
+Nonetheless, multiple studies indicate that graph transformers can effectively predict DDIs, outperforming existing algorithms [162].
 V-B2 Knowledge Graph Completion
-In the domain of Knowledge Graph (KG) completion, the utilization of graph transformers has been extensively investigated [174].
+In Knowledge Graph (KG) completion, graph transformers have been extensively studied [163].
 Chen et al.
-[175] proposed a novel inductive KG representation model, known as iHT, for KG completion through large-scale pre-training.
-This model comprises an entity encoder and a neighbor-aware relational scoring function, both parameterized by transformers.
-The application of this approach has led to remarkable advancements in performance, with a relative enhancement of more than 25% in mean reciprocal rank compared to previous models.
+[164] proposed iHT, an inductive KG representation model for KG completion through large-scale pre-training, featuring an entity encoder and a neighbor-aware relational scoring function, both transformer-parameterized.
+This approach achieved over 25% improvement in mean reciprocal rank compared to previous models.
 Liu et al.
-[176] introduced a generative transformer with knowledge-guided decoding for academic KG completion, which incorporates pertinent knowledge from the training corpus to provide guidance.
+[165] introduced a generative transformer with knowledge-guided decoding for academic KG completion, utilizing relevant knowledge from the training corpus.
 Chen et al.
-[177] developed a hybrid transformer with multi-level fusion to tackle challenges in multimodal KG completion tasks.
-This model integrates visual and textual representations through coarse-grained prefix-guided interaction and fine-grained correlation-aware fusion modules.
-V-B3 Recommender Systems
-Graph transformers have been effectively utilized in recommender systems by combining generative self-supervised learning with a graph transformer architecture [178].
+[166] developed a hybrid transformer with multi-level fusion for multimodal KG completion, integrating visual and textual representations via coarse-grained prefix-guided interaction and fine-grained correlation-aware fusion modules.
+V-B3 Recommendation Systems
+Graph transformers have been effectively employed in recommender systems by integrating generative self-supervised learning with a graph transformer architecture [167].
 Xia et al.
-[179] used the generative self-supervised learning method to extract representations from the data in an unsupervised manner and utilized graph transformer architecture to capture intricate relationships between users and items in the recommendation system.
+[168] utilized generative self-supervised learning to extract representations unsupervisedly and applied graph transformers to capture complex user-item relationships.
 Li et al.
-[180] introduced a new method for recommender systems that leverages graph transformers (GFormer).
-Their approach automates the self-supervision augmentation process through a technique called rationale-aware generative self-supervised learning.
-This technique identifies informative patterns in user-item interactions.
-The proposed recommender system utilizes a special type of collaborative rationale discovery to selectively augment the self-supervision while preserving the overall relationships between users and items.
-The rationale-aware self-supervised learning in the graph transformer enables graph collaborative filtering.
-While challenges remain in areas like graph construction, network design, model optimization, computation efficiency, and handling diverse user behaviors, experiments show that the approach consistently outperforms baseline models on various datasets.
+[169] introduced GFormer, which uses rationale-aware generative self-supervised learning to identify informative patterns in user-item interactions.
+This method automates self-supervision augmentation and preserves user-item relationships via collaborative rationale discovery.
+Despite challenges in graph construction, network design, model optimization, computation efficiency, and handling diverse user behaviors, experiments demonstrate that this approach consistently outperforms baseline models across various datasets [170].
 V-C Graph-level Tasks
@@
 V-C1 Molecular Property Prediction
-Graph transformers are powerful tools for molecular property prediction, utilizing the graph structure of molecules to capture essential structural and semantic information [181].
+Graph transformers are highly effective for molecular property prediction, leveraging molecular graph structures to capture essential structural and semantic information [171].
 Chen et al.
-[182] proposed Algebraic Graph-Assisted Bidirectional Transformer (AGBT) framework which integrates complementary 3D molecular information into graph invariants, rectifying the oversight of three-dimensional stereochemical information in certain machine learning models.
-Li et al [144] utilized Knowledge-Guided Pre-training of Graph Transformer (KPGT) in a self-supervised learning framework which emphasizes the importance of chemical bonds and models the structural information of molecular graphs.
+[172] introduced the Algebraic Graph-Assisted Bidirectional Transformer (AGBT) framework, which integrates 3D molecular information into graph invariants, addressing the lack of stereochemical information in some models.
+Li et al.
+[132] utilized Knowledge-Guided Pre-training of Graph Transformers (KPGT) in a self-supervised learning framework, emphasizing chemical bonds and modeling molecular graph structures.
 Buterez et al.
-[183] proposed transfer learning with graph transformer to enhance molecular property prediction on sparse and costly high-fidelity data.
+[173] proposed transfer learning with graph transformers to improve molecular property prediction on sparse, high-fidelity data.
 V-C2 Graph Clustering
-Graph transformers have been increasingly utilized in the field of Graph Clustering, offering innovative methodologies and overcoming significant challenges [184].
+Graph transformers are increasingly employed in Graph Clustering, offering innovative methods and addressing significant challenges [174].
 Yun et al.
-[185] proposed a graph transformer network to generate new graph structures for identifying useful connections between unconnected nodes on the original graph while learning effective node representation on the new graphs in an end-to-end fashion.
+[175] introduced a graph transformer network to generate new graph structures, identifying useful connections between previously unconnected nodes while learning effective node representations in an end-to-end manner.
 Gao et al.
-[186] proposed a patch graph transformer (PatchGT) that segments a graph into patches based on spectral clustering without any trainable parameters and allows the model to first use GNN layers to learn patch-level representations and then use transformer to obtain graph-level representations.
-These methodologies have addressed issues such as the limitations of the local attention mechanism and difficulties in learning high-level information, leading to enhanced graph representation, improved model performance, and effective node representation.
+[176] proposed the Patch Graph Transformer (PatchGT), which segments a graph into patches via spectral clustering without trainable parameters, using GNN layers to learn patch-level representations and transformers for graph-level representations.
+These methodologies overcome limitations of local attention mechanisms and difficulties in learning high-level information, resulting in enhanced graph representation, improved model performance, and effective node representation.
 V-C3 Graph Synthesis
-Graph transformers have been applied in the field of graph synthesis to improve graph data mining and representation learning.
-Existing graph transformers with Positional Encodings have limitations in node classification tasks on complex graphs, as they do not fully capture the local node properties [16].
+Graph transformers have advanced graph synthesis by enhancing graph data mining and representation learning.
+Traditional graph transformers with Positional encoding face limitations in node classification tasks on complex graphs due to inadequate capture of local node properties [16].
 To address this, Ma et al.
-[187] introduced the Adaptive Graph Transformer (AGT).
-This model tackles the challenge of extracting structural patterns from graphs in a way that is both effective and efficient.
-AGT achieves this by learning from two different graph perspectives: centrality and subgraph views.
-This approach has been shown to achieve state-of-the-art performance on real-world web graphs and synthetic graphs characterized by heterophily and noise.
+[177] introduced the Adaptive Graph Transformer (AGT), which effectively and efficiently extracts structural patterns from graphs by learning from centrality and subgraph views, achieving state-of-the-art performance on real-world web graphs and synthetic graphs with heterophily and noise.
 Jiang et al.
-[188] proposed an Anchor Graph Transformer (AGFormer) that leverages an anchor graph model to perform more efficient and robust node-to-node message passing for overcoming the computational cost and sensitivity to graph noises of regular graph transformers.
+[178] proposed the Anchor Graph Transformer (AGFormer), using an anchor graph model for efficient and robust node-to-node message-passing, overcoming computational costs and sensitivity to graph noise.
 Zhu et al.
-[122] developed a Hierarchical Scalable Graph Transformer (HSGT) which scales the transformer architecture to node representation learning tasks on large-scale graphs by utilizing graph hierarchies and sampling-based training methods.
+[114] developed the Hierarchical Scalable Graph Transformer (HSGT), scaling the transformer architecture to node representation learning tasks on large-scale graphs through graph hierarchies and sampling-based training methods.
 V-D Other Application Scenarios
@@
 V-D1 Text Summarization
-Text summarization is a crucial aspect of NLP which has been significantly advanced with the introduction of Graph transformers [190].
-These models utilize extractive, abstractive, and hybrid methods for summarization.
-Extractive summarization selects and extracts key sentences or phrases from the original text to create a summary [191].
-In contrast, abstractive summarization interprets the core concepts in the text and produces a concise summary.
-Hybrid summarization combines the advantages of both approaches [192].
-Despite the progress, challenges remain in text comprehension, main idea identification, and coherent summary generation.
-Nevertheless, the application of graph transformers in text summarization has demonstrated promising outcomes in terms of summary quality and efficiency [193].
+Text summarization, a crucial NLP task, has been significantly advanced by Graph transformers [180].
+These models employ extractive, abstractive, and hybrid methods for summarization [181].
+Extractive summarization selects key sentences or phrases from the original text [182], while abstractive summarization interprets core concepts to produce concise summaries.
+Hybrid summarization combines both approaches [183].
+Despite challenges in text comprehension, main idea identification, and coherent summary generation, graph transformers have shown promising results in summary quality and efficiency [184].
 V-D2 Image Captioning
-Graph transformers have emerged as a potent tool within the domain of image captioning, offering a structured representation of images and efficiently processing them to produce descriptive captions [194].
-Techniques such as Transforming Scene Graphs (TSG) leverage multi-head attention to architect graph neural networks for embedding scene graphs, which encapsulate a myriad of specific knowledge to facilitate the generation of words across various parts of speech [195].
-Despite encountering challenges, such as training complexity, absence of contextual information, and lack of fine-grained details in the extracted features, graph transformers have exhibited promising outcomes.
-They have enhanced the quality of generated sentences and attained state-of-the-art performance in image captioning endeavors [196].
+Graph transformers have become a powerful tool in image captioning, providing structured image representations and efficiently generating descriptive captions [185, 186].
+Techniques like Transforming Scene Graphs (TSG) use multi-head attention to design graph neural networks that embed scene graphs, capturing extensive knowledge to aid word generation across various parts of speech [187].
+Despite challenges like training complexity, lack of contextual information, and fine-grained details, graph transformers have shown promising results, improving sentence quality and achieving state-of-the-art performance in image captioning [188].
 V-D3 Image Generation
-Graph transformers have been effectively utilized in image generation, as demonstrated by various research studies [197].
+Graph transformers have been effectively utilized in image generation, as shown by various studies [189].
 Sortino et al.
-[33] proposed a transformer-based method conditioned by scene graphs for image generation, employing a decoder to sequentially compose images.
+[30] proposed a transformer-based method conditioned on scene graphs, employing a decoder for sequential image composition.
 Zhang et al.
-[198] proposed StyleSwin which uses transformers in constructing a generative adversarial network for high-resolution image creation.
-Despite challenges like redundant interactions and the requirement for intricate architectures, these studies have exhibited promising outcomes in terms of image quality and variety [199, 200].
+[190] introduced StyleSwin, using transformers in a generative adversarial network for high-resolution image creation.
+Despite challenges like redundant interactions and complex architectures, these studies have shown promising results in image quality and variety [120, 191].
 V-D4 Video Generation
-Graph transformers have been extensively applied in the field of Video Generation.
+Graph transformers have been extensively applied in video generation.
 Xiao et al.
-[201] proposed a Video Graph Transformer (VGT) model, which utilizes a dynamic graph transformer module to encode videos, capturing visual objects, their relationships, and dynamics.
-It incorporates disentangled video and text transformers for comparing relevance between the video and text.
+[192] introduced the Video Graph Transformer (VGT) model, using a dynamic graph transformer to encode videos by capturing visual objects, relationships, and dynamics, incorporating disentangled video and text transformers for relevance comparison.
 Wu et al.
-[202] proposed The Object-Centric Video Transformer (OCVT) which adopts an object-centric strategy to break down scenes into tokens suitable for a generative video transformer and understanding the intricate spatiotemporal dynamics of multiple interacting objects within a scene.
+[193] proposed the Object-Centric Video Transformer (OCVT), employing an object-centric approach to tokenize scenes for a generative video transformer, understanding complex spatiotemporal dynamics.
 Yan et al.
-[203] developed VideoGPT, which learns downsampled discrete latent representations of a raw video through 3D convolutions and axial self-attention.
-A GPT-like architecture is then used to model the discrete latent in a spatiotemporal manner using position encodings.
+[194] developed VideoGPT, which uses 3D convolutions and axial self-attention to learn downsampled discrete latent representations of raw videos, modeling these latents with a GPT-like architecture.
 Tulyakov et al.
-[204] proposed the MoCoGAN model which creates a video by mapping a series of random vectors to a sequence of video frames.
-Despite the challenges in capturing complex spatio-temporal dynamics in videos, these methodologies have exhibited promising outcomes across various facets of video generation, ranging from question answering to video summarization and beyond.
-VI Open Issues and Future Directions
+[195] proposed MoCoGAN, mapping random vectors to video frame sequences.
+Despite challenges in capturing complex spatiotemporal dynamics, these methodologies have shown promising outcomes in various video generation tasks, from question answering to video summarization.
+VI Design Guide for Graph Transformers
+Developing effective graph transformers requires meticulous attention to graph-specific challenges and careful consideration of design choices.
+This guide outlines practical principles for designing graph transformers across various graph types and tasks.
+  • Choose the appropriate type of graph transformer based on your graph data and tasks.
+Selecting a graph transformer requires a detailed understanding of the graph’s structure and scale.
+For ”simple, small” graphs, characterized by fewer than $10^{3}$ nodes and sparse connectivity (e.g., edge-to-node ratio <10<10), shallow transformers with fewer layers (e.g., 2–4 layers) may be sufficient to capture structural information efficiently.
+In contrast, ”complex, large” graphs, defined as having more than $10^{5}$ nodes, dense connectivity (edge-to-node ratio $\geq 10$), or rich attributes (e.g., temporal, weighted, or multi-relational edges), benefit from deeper models (e.g., 6–12 layers) capable of learning expressive representations.
+For dynamic or streaming graphs, where node or edge data evolves over time, scalable transformers leveraging techniques like subgraph sampling or incremental updates are essential [196].
+Sparse or noisy graphs, such as those with missing edges or incomplete attributes, are effectively handled by pre-trained graph transformers, which leverage knowledge transfer and task-specific fine-tuning to improve robustness.
+  • Design suitable structural and positional encodings for your graph data.
+Effective encoding of structural and positional information is crucial for transformers to capture graph-specific inductive biases.
+Techniques such as Laplacian eigenvector-based encodings, spectral embeddings, or random walk-based encodings (e.g., Node2Vec or DeepWalk embeddings) can encode global and local graph structure [197].
+For weighted graphs, edge weights can be integrated into positional encodings using distance-weighted features or flow-based encodings.
+For directed graphs, direction-sensitive encodings such as signed Laplacian eigenvectors or edge direction indicators can capture asymmetry [198].
+Incorporating these encodings into node or edge embeddings preserves structural informativeness and supports tasks such as graph classification and link prediction.
+  • Optimize the self-attention mechanism for your graph data.
+The self-attention mechanism in graph transformers computes pairwise attention scores, which are computationally expensive for large graphs [199].
+To address this, various optimization strategies can be employed:
+    • Sampling: Node sampling (e.g., GraphSAGE [84]) or edge sampling (e.g., DropEdge [200]) reduces computation by operating on representative graph subsets while preserving global properties.
+    • Sparsification: Techniques like top-$k$ sparsification retain only the most informative attention scores, reducing dense computations [201].
+    • Partitioning: Graph partitioning methods (e.g., METIS [202]) divide the graph into smaller, manageable subgraphs processed independently, improving scalability [203].
+    • Hashing: Methods such as locality-sensitive hashing approximate attention computations, enabling efficient processing of large node sets [204].
+    • Regularization and normalization: Regularizers like Dropout and attention normalization techniques mitigate overfitting and stabilize training, addressing issues like oversmoothing and oversquashing in deep transformers.
+    Each technique has trade-offs in terms of computational overhead, scalability, and performance fidelity, making the choice context-dependent.
+  • Utilize pre-training techniques to enhance performance.
+Pre-training techniques leverage large-scale datasets or pre-trained models to improve task-specific performance.
+Methods such as graph contrastive learning (e.g., GraphCL [205]) enhance feature extraction by maximizing agreement between augmented graph views.
+Fine-tuning techniques adapt pre-trained models for domain-specific tasks, while distillation transfers knowledge from large, resource-intensive models to smaller, efficient ones.
+Pre-training is particularly beneficial for low-resource scenarios, where labeled data is scarce, enabling effective downstream task performance with limited additional training.
+VII Open Issues and Future Directions
 Despite their immense potential for learning from graph-structured data, graph transformers still face open issues and challenges that require further exploration.
 Here we highlight some of these open challenges.
-VI-A Scalability and Efficiency
-The scalability and efficiency of graph transformers pose considerable challenges due to their substantial memory and computational requirements especially when employing global attention mechanisms to deal with large-scale graphs [205, 114].
-These challenges are further amplified in deep architectures which are susceptible to overfitting and over-smoothing.
-To address these issues, several potential strategies can be proposed:
+VII-A Scalability and Efficiency
+The scalability and efficiency of graph transformers face significant challenges due to high memory and computational demands, particularly with global attention mechanisms for large-scale graphs [206, 107].
+These challenges are exacerbated in deep architectures prone to overfitting and over-smoothing.
+To mitigate these issues, several strategies are proposed:
   1. 1.
-    Developing efficient attention mechanisms, such as linear, sparse and low-rank attention, to reduce the complexity and memory usage of graph transformers.
+    Developing efficient attention mechanisms, such as linear, sparse, and low-rank attention, to reduce complexity and memory usage.
   2. 2.
-    Applying graph sparsification or coarsening techniques to decrease the size and density of graphs while maintaining their crucial structural and semantic information.
+    Applying graph sparsification or coarsening techniques to decrease graph size and density while preserving key structural and semantic information.
   3. 3.
@@
   4. 4.
-    Exploring graph distillation or compression methods to create compact and effective graph transformer models, which are suitable for deployment on resource-limited devices.
+    Exploring graph distillation or compression methods to create compact and effective graph transformer models for deployment on resource-limited devices.
   5. 5.
-    Investigating regularization or normalization techniques, such as dropout, graph diffusion, convolution, and graph spectral normalization, to prevent overfitting and over-smoothing in graph transformer models.
-VI-B Generalization and Robustness
-Graph transformers often face challenges when it comes to generalizing to graphs that they have not encountered before or that fall outside of their usual distribution.
-This is especially true for graphs that have different sizes, structures, features, and domains.
-Additionally, graph transformers can be vulnerable to adversarial attacks and noisy inputs, which can result in a decline in performance and the production of misleading results.
-In order to improve the generalization and robustness of graph transformers, the following strategies could be taken into account:
+    Investigating regularization or normalization techniques, such as dropout, graph diffusion, convolution, and graph spectral normalization, to prevent overfitting and over-smoothing.
+VII-B Generalization and Robustness
+Graph transformers often struggle to generalize to unseen or out-of-distribution graphs, especially those with different sizes, structures, features, and domains.
+They are also vulnerable to adversarial attacks and noisy inputs, leading to performance drops and misleading results.
+To enhance the generalization and robustness of graph transformers, consider the following strategies:
   1. 1.
@@
   2. 2.
-    Applying domain adaptation or transfer learning techniques to facilitate learning from multiple source domains and transfer the knowledge from source domains to target domains.
+    Applying domain adaptation or transfer learning techniques to learn from multiple source domains and transfer knowledge to target domains.
   3. 3.
-    Exploring meta-learning or few-shot learning techniques to enable learning from limited data and rapid adaptation to new tasks.
+    Exploring meta-learning or few-shot learning techniques for learning from limited data and rapid adaptation to new tasks.
   4. 4.
@@
     Evaluating the uncertainty and reliability of graph transformer models using probabilistic or Bayesian methods, such as variational inference, Monte Carlo dropout, and deep ensembles.
-VI-C Interpretability and Explainability
-Graph transformers commonly regarded as black box models present significant challenges in terms of interpretability and explainability.
-The lack of sufficient justification and evidence for their decisions can undermine their credibility and transparency.
-To address this issue, several approaches can be considered:
+VII-C Interpretability and Explainability
+Graph transformers, often seen as black box models, face challenges in interpretability and explainability, undermining their credibility and transparency.
+To address this, several approaches can be considered:
   1. 1.
@@
   2. 2.
-    Applying explainable artificial intelligence (XAI) techniques, such as saliency maps, influence functions, and counterfactual explanations, to analyze and understand the behaviour and logic of graph transformer models.
+    Applying explainable artificial intelligence (XAI) techniques, such as saliency maps, influence functions, and counterfactual explanations, to analyze and understand the behavior and logic of graph transformer models.
   3. 3.
-    Exploring natural language generation techniques, such as template-based generation, neural text generation, and question-answering generation, to produce natural language explanations for outputs or actions of graph transformer models.
+    Exploring natural language generation techniques, such as template-based generation, neural text generation, and question-answering generation [207], to produce natural language explanations for outputs or actions of graph transformer models.
   4. 4.
     Investigating human-in-the-loop methods, such as active learning, interactive learning, and user studies, to incorporate human guidance in the learning or evaluation process of graph transformer models [28].
-VI-D Learning on Dynamic Graphs
-Graphs, which are frequently characterized by their dynamic and intricate nature, possess the ability to transform over time as a result of the addition or removal of nodes and edges, as well as the modification of node and edge attributes [206].
-Moreover, these graphs may exhibit diverse types and modalities of nodes and edges.
-In order to empower graph transformers to effectively manage such dynamic graphs, it is advisable to explore the following strategies:
+VII-D Learning on Dynamic Graphs
+Graphs are dynamic and complex, often changing over time with the addition or removal of nodes and edges, and the modification of their attributes [208, 209].
+They may also have diverse types and modalities of nodes and edges.
+To enable graph transformers to effectively handle such dynamic graphs, consider the following strategies:
   1. 1.
-    Developing temporal and causal attention mechanisms, such as recurrent, temporal, and causal attention, to capture the temporal and causal evolution of graphs.
+    Developing temporal and causal attention mechanisms, such as recurrent, temporal, and causal attention, to capture the temporal and causal evolution of graphs [210].
   2. 2.
-    Applying continual learning techniques on dynamic graphs to void forgetting previous knowledge and retraining.
+    Applying continual learning techniques on dynamic graphs to avoid forgetting previous knowledge and retraining.
   3. 3.
@@
     Leveraging multi-level and multi-layer attention mechanisms, such as node-edge, graph-graph, and hypergraph attention, to aggregate information from different levels and layers of graphs.
-VI-E Data Quality and Diversity
-Graph transformers require a significant quantity of diverse and high-quality data to achieve effective learning.
-Nevertheless, in the real world, data is frequently limited, noisy, incomplete, imbalanced, and biased.
-This has a detrimental effect on the performance and fairness of graph transformer models.
-To mitigate these challenges related to data quality and diversity, several potential strategies can be considered:
-  1. 1.
-    Developing noise-tolerant attention mechanisms, such as denoising attention, error correction attention, and self-supervised attention, to filter out data noise and errors.
-  2. 2.
-    Applying data augmentation or generation techniques, such as graph augmentation [207], graph completion, and graph generation, to enhance the quantity and variety of data.
-  3. 3.
-    Exploring data imbalance and bias mitigation techniques, such as data re-sampling, data re-weighting, and data debiasing, to improve data fairness and equity.
-  4. 4.
-    Evaluating data quality and diversity using data quality assessment or data diversity analysis methods, such as data quality metrics, data diversity measures, and data quality visualization.
-VII Conclusion
+VIII Conclusion
 Graph transformers are a novel and powerful class of neural network models, which can effectively encode and process graph-structured data.
 This survey provides a comprehensive overview of graph transformers in terms of design perspectives, taxonomy, applications, and open issues.
-We first discuss how graph transformers incorporate graph inductive bias, including node positional encodings, edge structural encodings, message-passing bias and attention bias, to encode the structural information of graphs.
+We first discuss how graph transformers incorporate graph inductive bias, including node positional encoding, edge structural encoding, message-passing bias and attention bias, to encode the structural information of graphs.
 Then, we introduce the design of graph attention mechanisms, including global and local attention mechanisms.
@@
 This survey aims to serve as a valuable reference for researchers and practitioners interested in graph transformers and their applications.
-Acknowledgment
-The authors would like to thank Estrid He (RMIT University) for her helpful comments.